### v2

# The gauge symmetry - unbroken and broken - in the QCD ground state

Abstract
The main purpose of this work is to report on convincing progress made in the investigation of the one of the long-standing conceptual problems in QCD, namely whether the gauge symmetries of the QCD Lagrangian and its ground state coincide or not? For this purpose, we use only the tensor algebra derivation rigorous rules in dealing with the corresponding skeleton loop integrals, contributing into the full gluon self-energy. It has been exactly (i.e., without making any kind of the simplifications) proven that in the QCD ground state may exist the two principally different types of the gluon field configurations. One of them maintains the QCD Lagrangian’s gauge symmetry, but another one does not.The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared. The exact proof has been given that it cannot be disregarded from the theory and its ground state by any means. The gauge symmetry of the QCD Lagrangian forbids such term, but the perturbative renormalizability of the theory will not be affected. This allows to present the new significant insights into the gauge and dynamical structures of the true QCD ground state, and sheds light on a mass squared dynamical generation in it.
keywords:
theoretical particle physics, Quantum Chromodynamics, ground state, gluon full self-energy, dynamical and gauge structures.
PACS:
11.10.-z, 11.15.-q, 12.38.-t, 12.38.Aw, 12.38.Lg
MSC:
[2020] 30D30, 81Q40, 81T13, 81T16
1 Introduction
The quark model (QM) treats the strongly-interacting hadrons (baryons and mesons) as bound-states of quarks, by emitting and absorbing gluons. The theory, which purpose is to describe the properties of the observed hadrons in terms of the non-observable quarks and gluons from first principles, is Quantum Chromodynamics (QCD) 1 ; 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 . It is widely accepted as the quantum gauge field theory of strong interactions. Being such a theory, it has to undergo the two phase transitions: the first one at the fundamental (microscopic) quark-gluon level is – the confinement phase transition – in order to explain why all the physical states are colour-singlets. The second one at the (macroscopic) hadronic level is – the PCAC (partially conserved axial currents) phase transition – in order to explain the soft pion physics 2 ; 23 .
However, this purpose remains a formidable task yet because of the multiple dynamical and topological complexities of low-energy particles (hadrons) physics, originated from QCD and its ground state (vacuum). This happens because QCD as a gauge theory of the gluon-gluon and quark-gluon strong interactions suffers from a few important conceptual problems. Here we focus on the two of them, being in the close connection with each other.
A. Whether the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is a symmetry of its ground state or not?
B. The dynamical generation of a mass squared in the vacuum of QCD, since the gauge invariance of its Lagrangian forbids such kind of terms, like the gluon mass term $M^{2}_{g}A_{\mu}A_{\mu}$.
The solutions of these fundamental problems is a key to the explanation of the first confinement phase transition in QCD. The properties and symmetries of the QCD Lagrangian, and thus including its Yang – Mills (YM) part, are well-known 1 ; 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 ; 10 . The propagation of gluons is one of the main dynamical effects in the QCD ground state. The importance of the corresponding equation of motion is due to the fact that its investigation is supposed to reflect the dynamical and gauge structures of the QCD ground state. The gluon Schwinger – Dyson (SD) equation 2 ; 10 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 is highly non-linear (NL) because of the self-interaction of massless gluon modes. They are taken into account by the full gluon self-energy, which is the main part of the gluon SD equation itself. We will show that the general properties of the full gluon self-energy point out on some new dynamical and gauge aspects of the true QCD ground state. Due to these inputs, the novel insights into a mass dynamical generation in the QCD vacuum are also present within proposed below formalism, based on the tensor algebra derivation rigorous rules only.
2 The gluon SD equation
The structure of the gluon SD equation is given in this section in some necessary details. Analytically it looks like
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{0}_{\mu\rho}(q)i\Pi_{\rho\sigma}(q;D)D_{\sigma\nu}(q),$ (2.1)
where $D_{\mu\nu}(q)$ and $D^{0}_{\mu\nu}(q)$ denote the full gluon propagator and its free counterpart, respectively. $\Pi_{\rho\sigma}(q;D)$ is the full gluon self-energy which depends on the full gluon propagator due to the non-abeian character of QCD. Here and everywhere below we omit the colour group indices, for simplicity, because of their final factorization, for example $D^{ab}_{\mu\nu}(q)=D_{\mu\nu}(q)\delta^{ab}$. Eq. (2.1) in terms of the corresponding skeleton loop diagrams is shown in Fig. 1.
Figure 1: The SD equation for the full gluon propagator as present in 10 .
Here stringy lines are for the free gluon propagator, while $D$ denotes its full counterpart. $S$ with solid lines denotes the full quark propagator, and $\Gamma$ denotes the full quark-gluon vertex. $G$ with dashed lines denotes the full ghost propagator, and $G_{\mu}$ is the full ghost-gluon vertex. Finally, $T_{3}$ and $T_{4}$ denote the full 3- and 4-gluon vertices, respectively.
The full gluon self-energy is convenient to present as the sum of the three independent terms, namely
$\Pi_{\rho\sigma}(q;D)=\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)+\Pi_{\rho\sigma}^{t}(D),$ (2.2)
where $\Pi^{q}_{\rho\sigma}(q)$ describes the skeleton loop contribution of the quark degrees of freedom as an analogue to the vacuum polarization tensor in Quantum Electrodynamics (QED) 7 ; 21 . Note that here and below the superscript or subscript ’$q$’ means quark (not to be mixed up with the gluon momentum variable $q$). The gluon part of the full gluon self-energy by itself is the sum of a few independent terms as follows:
$\Pi^{g}_{\rho\sigma}(q;D)=\Pi^{gh}_{\rho\sigma}(q)+\Pi^{(1)}_{\rho\sigma}(q;D^{2})+\Pi^{(2)}_{\rho\sigma}(q;D^{4})+\Pi^{(2^{\prime})}_{\rho\sigma}(q;D^{3}),$ (2.3)
and $\Pi^{gh}_{\rho\sigma}(q)$ describes the skeleton loop contribution associated with the ghost degrees of freedom. $\Pi^{(1)}_{\rho\sigma}(q;D^{2})$ represents the skeleton loop contribution, containing the 3-gluon vertices only. Finally, $\Pi^{(2)}_{\rho\sigma}(q;D^{4})$ and $\Pi^{(2^{\prime})}_{\rho\sigma}(q;D^{3})$ describe the skeleton two-loop contributions, which combine the 3- and 4-gluon vertices. All these quantities are given by the corresponding skeleton loop diagrams in Fig. 1, and they are independent from each other. The last four terms explicitly contain the full gluon propagators in the corresponding powers symbolically shown above. They form the NL part of the gluon SD equation. The analytical expressions for the corresponding skeleton loop integrals 22 , in which the symmetry and combinatorial coefficients and signs have been included, are not important here. We are not going to calculate any of them explicitly, and thus to introduce into them any truncations/approximations/assumptions or choose some special gauge. Any skeleton loop integral in Fig. 1 is the sum of the infinite number of terms. Moreover, the full vertices entering these skeleton loop integrals are themselves determined by an infinite series of the corresponding multi-loop skeleton terms 2 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 . In these NL series the dependence on the coupling constant may not be simple, i.e., in fact, it is not known. Such kind of series are the so-called ’cluster’ expansions 23 , indeed.
The constant tadpole term $\Pi_{\rho\sigma}^{t}(D)$ is defined as follows:
$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lD_{\alpha\beta}(l)T^{0}_{\rho\sigma\alpha\beta}=g_{\rho\sigma}\Delta^{2}_{t}(D)=[T_{\rho\sigma}(q)+L_{\rho\sigma}(q)]\Delta^{2}_{t}(D),$ (2.4)
where $L_{\rho\sigma}(q)=q_{\rho}q_{\sigma}/q^{2}$, so that the tadpole term contributes into the both transverse and longitudinal components of the full gluon propagator (2.1). In what follows the constant $\Delta^{2}_{t}(D)$ will be called as the tadpole term itself, since it deviates from the definition (2.4) by the tensor $g_{\rho\sigma}$ only.
It is instructive for further purpose to present the initial gluon SD eq. (2.1), on account of the expressions (2.2) and (2.4) as follows:
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)$ $+$ $D^{0}_{\mu\rho}(q)i\left[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)\right]D_{\sigma\nu}(q)$ (2.5)
$+$ $D^{0}_{\mu\rho}(q)ig_{\rho\sigma}\Delta^{2}_{t}(D)D_{\sigma\nu}(q).$
All the terms which contribute to the full gluon self-energy eq. (2.2), and hence eq. (2.5), are tensors, having the dimensions of a mass squared. All these skeleton loop integrals are therefore quadratically divergent (QD), i.e., ultraviolet (UV) divergent in the perturbation theory (PT) regime, and so they are assumed to be regularized. Contrary to QED, QCD being a non-abelian gauge theory can suffer from the severe infrared (IR) singularities in the $q^{2}\rightarrow 0$ limit (i.e., more singular than the PT one, which is proportional to $\sim 1/q^{2}$). Evidently, they will be generated due to the self-interaction of massless gluon modes. Thus, all the possible subtractions at zero may be dangerous 2 ; 3 ; 4 ; 5 ; 6 ; 7 . That is why in all the quantities below, the dependence on the finite (slightly different from zero) dimensionless subtraction point $\alpha$ has to be understood. In other words, all the subtractions at zero and the Taylor expansions around zero should be understood as the subtractions at $\alpha$ and the structure of the Taylor expansions near $\alpha$, where they are justified to be used. From the technical point of view, however, it is convenient to put formally $\alpha=0$ in all the expressions and derivations below, and to restore the explicit dependence on non-zero $\alpha$ in all the quantities only at the later stage. In all the quantities where the dependence on the dimensionless UV regulating parameter $\tilde{\lambda}$ and $\alpha$ is not shown explicitly, nevertheless, it should be also assumed. For example, $\Pi_{\rho\sigma}(q;D)\equiv\Pi_{\rho\sigma}(q;D,\tilde{\lambda},\alpha)$ and similarly for all other quantities. This means that all the expressions are regularized (i.e., we render infinite skeleton loop integrals to be finite).
The tadpole term $\Delta^{2}_{t}(D)$, defined in eq. (2.4), is nothing else but the QD skeleton loop integral, which is already regularized from below and above as well as all the other such kind of constants which may appear in the theory, as a results of the proper subtraction scheme (see below). Being a mass squared regularized quantity, it is explicitly present in the QCD ground state, see eq. (2.5). However, it violets the gauge symmetry of the QCD Lagrangian. Then a natural question arises why is it present in the vacuum of QCD at all if it makes theory to look un- renormalizable from the very beginning? The standard solution to this dilemma was to remove this and all other QD constants in order to make theory renormalizable in accordance with the QCD Lagrangian’s exact gauge symmetry. However, here we are going to solve this dilemma by asking the question is it possible to retain the tadpole term in the QCD vacuum, but without affecting the PT renormalization properties of the theory? This is important to ask because, unlike to other QD constants, the tadpole term is only one which generates a mass squared scale parameter 6 . Also, one of our primary aims will be to clarify its role in the dynamical and gauge structures of the QCD vacuum in more detail. For this purpose we will develop a suitable formalism, based only on the rigorous rules of the tensor algebra derivations.
Concluding, let us remind that the whole gluon momentum range is $q^{2}\in[0,\infty)$. In what follows we will work in Euclidean metric $q^{2}=q^{2}_{0}+\bf q^{2}$ since it implies $q_{i}\rightarrow 0$ when $q^{2}\rightarrow 0$ and vice versa. This makes it possible to avoid the unphysical IR singularities on the light cone.
3 Slavnov-Taylor identities
The basic relation to which the full gluon propagator should satisfy is the corresponding Slavnov-Taylor (ST) identity 2 ; 24 ; 25 , namely
$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi,$ (3.1)
where $\xi$ is the gauge-fixing parameter. It implies that the general tensor decomposition of the full gluon propagator in the covariant gauge is as follows:
$D_{\mu\nu}(q)=i\left[T_{\mu\nu}(q)d(q^{2})+\xi L_{\mu\nu}(q)\right]{1\over q^{2}},$ (3.2)
where the invariant function $d(q^{2})$ is the corresponding Lorentz structure of the full gluon propagator (or the gluon invariant function). Throughout this paper we use the standard definition of $T_{\mu\nu}(q)=\delta_{\mu\nu}-q_{\mu}q_{\nu}/q^{2}=\delta_{\mu\nu}-L_{\mu\nu}(q)$ in Euclidean metric. Any invariant functions associated with the projective operators $T_{\mu\nu}(q)$ and $\delta_{\mu\nu}$ are the same, and thus $D_{\mu\nu}(q)$ is defined up to its longitudinal part $L_{\mu\nu}(q)$. This should be also valid for its equation of motion.
If one neglects all the contributions to the full gluon self-energy in eq. (2.1), by setting $d(q^{2})=1$ in eq. (3.2), then one obtains the free gluon propagator
$D^{0}_{\mu\nu}(q)=i\left[T_{\mu\nu}(q)+\xi_{0}L_{\mu\nu}(q)\right](1/q^{2}),$ (3.3)
where $\xi_{0}$ is the corresponding gauge-fixing parameter. The general ST identity (3.1) will look like
$q_{\mu}q_{\nu}D^{0}_{\mu\nu}(q)=i\xi_{0}.$ (3.4)
Let us note in advance that from the very beginning the gauge-fixing parameter for the full gluon propagator $\xi$ is not to be identified with its free counterpart $\xi_{0}$. Under which conditions they are the same or different is one of the main subjects of our investigation in this paper. Since the gauge freedom in QCD is determined by $\xi_{0}$, the gauge-fixing parameter $\xi$ has to be a function of $\xi_{0}$ i.e., $\xi=f(q^{2};\xi_{0})$ and called the gauge-fixing function.
Emphasizing, these and all other ST identities which appear in QCD 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 9 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 are important for the general renormalizability of the theory, and thus they ”are exact constraints on any solution to QCD” 2 .
3.1 Transverse relations for the full gluon self-energy
Let us now perform some algebraic derivations which are necessary for the further purposes. Contracting the full gluon SD eq. (2.1) and eq. (2.5) with $q_{\mu}$ and $q_{\nu}$, on account of the relations (3.1)-(3.4), one gets
$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)$ $=$ ${(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2},$
$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]$ $=$ ${(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}-q^{2}\Delta^{2}_{t}(D),$ (3.5)
and they related to each other through the relation (2.2) as well as eq. (2.1) is related to eq. (2.5), which is as it should be. The first line in the transverse relations (3.5) is the tensor relation, while its second one is nothing else but the tensor identity. The second transverse relation explicitly depends on the tadpole term $\Delta^{2}_{t}(D)$. If it is formally put zero, then the both transverse relations (3.5) coincide with each other. This underlines its important role in the QCD ground state, which one precisely is to be clarified in our paper.
[…]
However, by this mechanism the QD constants due to the quark degrees of freedom only can be excluded from the QCD theory. What is about the exclusion of the QD constants from its YM sector? QCD being much more complicated quantum field gauge theory than QED, has the two transverse relations (3.5), second of which shows the gluon degrees of freedom explicitly. By solution of these relations, we mean that at least one of them has to be necessarily satisfied. Finally, this will lead to the removal of the corresponding QD constants from the YM theory as well as to fix the function $\xi=f(q^{2};\xi_{0})$, see exact derivations in sections 4, 5 and 6. This strategy is definitely accessible with conventional means of the gauge theories to make them renormalizable, and QCD is a such theory 2 , indeed. So that whether these transverse relations will be satisfied/saturated (i.e., equal zero) or not and further include them into the proper subtraction scheme requires much more careful investigation in QCD. Obviously, there exist only the two different requirements in order to satisfy at least one of the transverse relations (3.5). However, it is time now to introduce the proper subtraction scheme.
4 The proper subtraction scheme
The first step in the renormalization program of any gauge theory is the removal of the quadratic UV divergences in order to make the corresponding theory renormalizable in the PT sense. It can be achieved by introducing the proper subtraction scheme in order to separate them from the PT logarithmic divergences. The preliminary step in the regularization program has been already done by introducing the corresponding regulating parameters $\tilde{\lambda}$, $\alpha$.
Within our approach nothing will depend on how exactly the regulating parameters have been introduced. Even the hard cut-off procedures can be used, since we are mainly interested in the regulating of leading order divergences 2 . They have to disappear from the theory after the PT and the non-perturbative (NP) renormalization programs will be performed.
Let us now introduce the subtractions for the quark and gluon contributions to the full gluon self-energy as follows:
$\Pi^{s(q)}_{\rho\sigma}(q)$ $=$ $\Pi^{q}_{\rho\sigma}(q)-\Pi^{q}_{\rho\sigma}(0)=\Pi^{q}_{\rho\sigma}(q)-\delta_{\rho\sigma}\Delta^{2}_{q},$
$\Pi^{s(g)}_{\rho\sigma}(q;D)$ $=$ $\Pi^{g}_{\rho\sigma}(q;D)-\Pi^{g}_{\rho\sigma}(0;D)=\Pi^{g}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{g}(D),$ (4.1)
and thus $\Pi^{s(q)}_{\rho\sigma}(0)=\Pi^{s(g)}_{\rho\sigma}(0;D)=0$, by definitions. In these relations $\Delta^{2}_{q}$ is the skeleton loop integral for the quark degrees of freedom at $q=0$, and $\Delta^{2}_{g}(D)$ is the sum of the corresponding skeleton loop integrals at $q=0$ contributing to eq. (2.3), namely
$\Delta^{2}_{g}(D)=\Delta^{2}_{gh}+\Delta^{2}_{1}(D^{2})+\Delta^{2}_{2}(D^{4})+\Delta^{2}_{2^{\prime}}(D^{3}).$ (4.2)
Let us remind that all the QD constants, shown up in eqs. (4.1)-(4.2), as well as the tadpole term $\Delta^{2}_{t}(D)$ itself, are independent from each other and are defined by the skeleton loop integrals, which have been already regularized from above and below. The subtraction at zero is to be understood in a such way that we subtract at the external gluon momentum $q^{2}=-\mu^{2}$ 2 with $\mu^{2}\rightarrow 0$ final limit (its dimensionless counterpart is $\alpha$ mentioned above). However, due to the self-interaction of the massless gluon modes, some of these constants (apart from the tadpole term) may depend on the internal zero gluon momentum (if it is connected to any closed loop, for example see appendix A). Then it has to be replaced by the subtraction at $q^{2}_{i}=-M^{2}_{i}$ with $M^{2}_{i}\rightarrow 0$ final limit as well, and subindex $i$ determines the number of such internal gluons. For the any full gluon propagator the subindex $i$ can be treated as the number of the necessary subtractions made in it in accordance with the rules of the theory of distributions (generalized functions) 26 .
Also it is necessary to emphasize that the subtractions (4.1) have not been introduced by hand. They are equivalent to add zero to the corresponding identity. For example $\Pi^{q}_{\rho\sigma}(q)=\Pi^{q}_{\rho\sigma}(q)-\Pi^{q}_{\rho\sigma}(0)+\Pi^{q}_{\rho\sigma}(0)$, so that the initial quark contribution to the gluon self-energy is not changed. It makes it possible to present the initial quark contribution $\Pi^{q}_{\rho\sigma}(q)=\Pi^{s(q)}_{\rho\sigma}(q)+\Pi^{q}_{\rho\sigma}(0)$ as a sum of the two terms, one of which shows up all the corresponding QD but regularized constants. The first term is a regular function of the external gluon momentum $q$, by definition, and it may be only logarithmically divergent at large $q$. The same is true for the initial gluon contribution as well. Such an exact separation will be very useful for the renormalization programs of any kind to be performed for the single full gluon propagator, and we consider this as an advantage of the formalism developed here. For the interacting full gluon propagators these independent terms will interact with each other, but such a separation will simplify the corresponding renormalization programs, anyway.
The independent tensor decompositions of the quark and gluon degrees of freedom, which appear in the subtraction relations (4.1), are
$\Pi^{s(q)}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(q)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(q)}_{l}(q^{2}),$
$\Pi^{q}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{q}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{q}_{l}(q^{2}),$ (4.3)
and
$\Pi^{s(g)}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(g)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(g)}_{l}(q^{2}),$
$\Pi^{g}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{g}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{g}_{l}(q^{2}),$ (4.4)
respectively. In all the quantities above and below the dependence on $D$ is omitted, for simplicity, and will be restored when necessary. Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary. However, all the invariant functions $\Pi^{s(q)}_{t}(q^{2})$, $\Pi^{s(q)}_{l}(q^{2})$ as well as $\Pi^{s(g)}_{t}(q^{2})$, $\Pi^{s(g)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the subtractions (4.1).
Substituting all these decompositions (4.3)-(4.4) into the subtractions (4.1), and doing some tensor algebra derivations, one finally obtains
$\Pi^{s(q)}_{t}(q^{2})=\Pi^{q}_{t}(q^{2})-{\Delta^{2}_{q}\over q^{2}},\quad\Pi^{s(q)}_{l}(q^{2})=\Pi^{q}_{l}(q^{2})+{\Delta^{2}_{q}\over q^{2}},$ (4.5)
and
$\Pi^{s(g)}_{t}(q^{2})=\Pi^{g}_{t}(q^{2})-{\Delta^{2}_{g}(D)\over q^{2}},\quad\Pi^{s(g)}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})+{\Delta^{2}_{g}(D)\over q^{2}}.$ (4.6)
Using these relations it is easy to show that for the regularized (i.e., finite) quark and gluon contributions to the initial gluon self-energy their values at zero always are $\Pi^{q}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{q}$ and $\Pi^{g}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{g}(D)$, indeed, i.e, they are general ones, and have been already used in the subtraction relations (4.1) above, as it has to be. Also these relations will be true if expressed in terms of the above-mentioned subtraction point $q^{2}=-\mu^{2}$.
Let us now make an important observation in advance. If the initial contributions of the quarks and gluons degrees of freedom to the full gluon self-energy, which are present by the second lines in the relations (4.3) and (4.4), respectively, are not transverse, then one cannot remove the corresponding QD constants $\Delta^{2}_{q}$ and $\Delta^{2}_{g}(D)$ from the theory. On the contrary, if their are transverse, i.e., $\Pi^{q}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})=0$, then from the relations (4.5) and (4.6) it follows that one has to put the corresponding QD constants $\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0$, and thus removing them from the theory, since the corresponding subtracted invariant functions cannot have the pole-type singularities, by definition. This observation explicitly shows the necessity to satisfy one of the transverse relations (3.5) from the general point of view.
An interesting observation follows from the relations (4.5) and (4.6) as well, namely $\Pi^{q}_{t}(q^{2})+\Pi^{q}_{l}(q^{2})=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(q)}_{l}(q^{2})$ and $\Pi^{g}_{t}(q^{2})+\Pi^{g}_{l}(q^{2})=\Pi^{s(g)}_{t}(q^{2})+\Pi^{s(g)}_{l}(q^{2})$. So that the corresponding sums do not depend on the corresponding QD but regularized constants, which once more underlines the general character of our subtraction scheme. Also, the sum of the subtracted invariant functions is never zero, though the subtracted contributions to the full gluon self-energy are zero at $q^{2}=0$, by definitions, see relations (4.1).
5 Gauge symmetry is unbroken
In order to satisfy the first transverse relation (3.5) one needs to put $\xi=f(\xi_{0})=\xi_{0}$, and thus it becomes transverse $q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)=0$, indeed. Then from the second of the relations (3.5) it follows that
$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]=-q^{2}\Delta^{2}_{t}(D),$ (5.1)
and it is not transverse, as expected. Substituting further the subtractions (4.1) into the relation (5.1), and doing some algebra, one obtains
$q_{\rho}q_{\sigma}[\Pi^{s(q)}_{\rho\sigma}(q)+\Pi^{s(g)}_{\rho\sigma}(q;D)]=-q^{2}\Delta^{2}(D),$ (5.2)
where
$\Delta^{2}(D)=\Delta^{2}_{q}+\Delta^{2}_{g}(D)+\Delta^{2}_{t}(D),$ (5.3)
is the sum of all the QD but regularized constants, which may appear in the theory.The QD but regularized constant $\Delta^{2}_{g}(D)$ itself is defined in eq. (4.2).
Substituting further the first tensor decompositions of the relations (4.3) and (4.4) into the transverse relation (5.2), and doing some algebra, one obtains for their longitudinal invariant functions the following relation
$\Pi^{s(q)}_{l}(q^{2})+\Pi^{s(g)}_{l}(q^{2})={\Delta^{2}(D)\over q^{2}},$ (5.4)
which is impossible, since these invariant functions cannot have the pole-type singularities, by definition, as underlined just above. So that the QD but regularized constant $\Delta^{2}(D)$ should be discarded from the theory, i.e., put formally zero $\Delta^{2}(D)=0$ on this mathematical basis. Then $\Pi^{s(q)}_{l}(q^{2})=\Pi^{s(g)}_{l}(q^{2})=0$ as well, since they are independent from each other.
From now on in this section, we replace $D$ by $D^{PT}$ (for the explanation see the text below). The definitions (4.2) and (5.3) require that this system of the relations for the QD but regularized constants now looks like
$\Delta^{2}_{q}=\Delta^{2}_{g}(D^{PT})=\Delta^{2}_{t}(D^{PT})=0,$ (5.5)
and
$\Delta^{2}_{gh}=\Delta^{2}_{1}(D^{2}_{PT})=\Delta^{2}_{2}(D^{4}_{PT})=\Delta^{2}_{2^{\prime}}(D^{3}_{PT})=0,$ (5.6)
because all of them are independent from each other. The dimensional regularization method (DRM) 4 ; 6 ; 27 ; 28 ; 29 provides a gauge-invariant scheme to correctly calculate the finite parts of the generally QD loop integrals, while omitting their QD parts, i.e., simply to ignore them and to deal further only with the logarithmic divergences of the PT. This was a prescription rather than an exact result, as pointed out in 4 . But the relations (5.5)-(5.6) are not prescriptions, they are mathematically exact results. How to understand such equalities is convenient to postpone until next section.
Collecting now all our results obtained above, one arrives at the system of the relations as follows:
$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D^{PT})=0,$ (5.7)
and
$\xi=f(\xi_{0})=\xi_{0},\quad\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D^{PT})=T_{\rho\sigma}(q)q^{2}\Pi^{s}(q^{2};D^{PT}),$ (5.8)
where $\Pi^{s}(q^{2};D^{PT})=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(g)}_{t}(q^{2};D^{PT})$, since the quark and gluon contributions coincide with their subtracted counterparts and are transverse in this case as well, i.e., $\Pi^{q}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})=0$, due to the second equations in the relations (4.5)-(4.6). The invariant function $\Pi^{s}(q^{2};D^{PT})$ is a regular function at small $q^{2}$ and may be only logarithmically divergent at large $q^{2}$.
The gluon SD eq. (2.5) becomes
$D^{PT}_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)q^{2}\Pi^{s}(q^{2};D^{PT})D^{PT}_{\sigma\nu}(q),$ (5.9)
and from it one arrives at $q_{\mu}q_{\nu}(q)D^{PT}_{\mu\nu}(q)=q_{\mu}q_{\nu}D^{0}_{\mu\nu}(q)=i\xi_{0}$, and thus the gauge-fixing parameter of the PT full gluon propagator is fixed as follows: $\xi=\xi_{0}$, in complete agreement with eqs. (5.8). Substituting decompositions (3.2) and (3.3) into the eq. (5.9), one finally obtains
$D^{PT}_{\mu\nu}(q)={iT_{\mu\nu}(q)\over q^{2}[1+\Pi^{s}(q^{2};D^{PT})]}+i\xi_{0}L_{\mu\nu}(q){1\over q^{2}},$ (5.10)
so that the gluon invariant function is $d_{PT}(q^{2})=[1+\Pi^{s}(q^{2};D^{PT})]^{-1}$. Obviously, this equation describes the propagation of the PT massless gluons, since it has the PT singularity on the mass-shell $q^{2}=0$, i.e., the singularity of the free gluon propagator $\sim 1/q^{2}$. Let us note in advance that the equality $\xi=\xi_{0}$ takes place only for the regularized massless gluon fields. For their renormalized counterparts these gauge-fixing parameters differ by the corresponding renormalization constant 2 ; 3 ; 4 ; 5 ; 6 ; 7 .
The absence of a mass scale parameters in eq. (5.10) can be now attributed to the satisfied transverse relations (5.7)-(5.8). It is possible to say that just these ones decrease the quadratic UV divergences of the corresponding skeleton loop integrals to a logarithmic ones at large $q^{2}$. Due to the transverse relations (5.7)-(5.8) all the scale parameters, having the dimensions of a mass squared, shown up in eqs. (5.5)-(5.6), should be removed/disappeared from the theory, i.e., put formally zero.
The system of the regularized eqs.(5.9)-(5.10), which is uniquely defined, is free of all the types of the scale parameters having the dimensions of a mass squared, forbidden by the exact gauge symmetry of the QCD Lagrangian. Therefore, it constitutes that the gauge symmetry of the QCD ground state, reflected by this system, coincides with the symmetry of its Lagrangian. Such coincidence is analogous to QED where the abelian $U(1)$ gauge symmetry of the Lagrangian is the same as of its ground state (omitting formally the YM degrees of freedom here, this coincidence has been proven just in this way in 10 ). However, it is necessary to remind that the coupling constant remains strong, so to call it as $D^{PT}$ is only a mere convention.
The essential feature of the above-mentioned coincidence is the equality $\xi=\xi_{0}$. In this connection, let us note that one can start from the expression (3.2). It is the general decomposition of the full gluon propagator, depending on the one variable, into the independent tensor structures. Then it will automatically satisfy to the relation (3.1), called the ST identity, determining thus the gauge-fixing parameter. This has nothing to do with the statement that the ST identity is a consequence of the exact gauge symmetry. It becomes its consequence, indeed, when $\xi$ is fixed by the relation (5.8). The well-known PT linear relation $\xi=\xi_{0}$ is the consequence of the exact gauge invariance both for the QCD Lagrangian and its ground state if it is treated in the PT framework for the regularized gluon fields. In principle, the PT QCD is a well-elaborated quantum gauge field theory 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 , and it is not our concern. We present it here within our formalism for the readers convenience to directly compare it with our approach to the QCD below.
6 Gauge symmetry is broken!
QCD, being a non-abelian gauge theory with the strong coupling constant, is to be treated beyond the PT, especially its ground state. It contains much more information on the dynamical and gauge structures of the theory than its Lagrangian can provide at all 2 . So that the general case when $\xi\neq\xi_{0}$ should be investigated in detail, while keeping the ST identities (3.1) and (3.4) valid, anyway. The another independent requirement to satisfy the second one of the transverse relations (3.5) is defined by the condition
${(\xi_{0}-\xi)\over\xi\xi_{0}}q^{2}-\Delta^{2}_{t}(D)=0.$ (6.1)
Then for the initial gluon self-energy its transverse relation looks like
$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)=q^{2}\Delta^{2}_{t}(D),$ (6.2)
and it is not transverse, while the second line of the system (3.5) becomes transverse,
$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]=0,$ (6.3)
as expected and compare it with eq. (5.1). In the previous section we do not use that the contribution of the quark skeleton loop to the full gluon self-energy can be made always transverse, i.e., put $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=0$, independently from its gluon counterpart (this is an analogue to the satisfied transverse relation for the vacuum polarization tensor in QED, mentioned above). So that the satisfied transverse relation (6.3) can be effectively splintered into the two independent transverse relations, namely
$q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=0,\quad q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0.$ (6.4)
However, from the second lines of the tensor decompositions (4.3)-(4.4) and the satisfied transverse relations (6.4), one arrives to the conditions $\Pi^{q}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})=0$, like in the previous case. Then the second relations in eqs (4.5) and (4.6) are to be reduced to
$\Pi^{s(q)}_{l}(q^{2})={\Delta^{2}_{q}\over q^{2}},\quad\Pi^{s(g)}_{l}(q^{2};D)={\Delta^{2}_{g}(D)\over q^{2}},$ (6.5)
but that is impossible, since the corresponding invariant functions cannot have pole-type singularities, by definitions, as explained in the previous section. One has to put these constants to zero on this general mathematical basis, i.e., $\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0$, and $\Pi^{s(q)}_{l}(q^{2})=\Pi^{s(g)}_{l}(q^{2};D)=0$ as well.
All our results for the QD constants can be now summarized as follows:
$\Delta_{q}^{2}=\Delta_{g}^{2}(D)=0,\quad\Delta^{2}(D)\rightarrow\Delta_{t}^{2}(D)\neq 0,$ (6.6)
and hence
$\Delta^{2}_{gh}=\Delta^{2}_{1}(D^{2})=\Delta^{2}_{2}(D^{4})=\Delta^{2}_{2^{\prime}}(D^{3})=0,$ (6.7)
as well because of the relation (4.2). It is instructive to compare this system of the relations with the relations (5.5)-(5.6). The common feature of these relations is that the quark and gluon constants should be removed from the theory not only in the PT QCD but in the QCD as well. This means that the correspondence with the PT system (5.5)-(5.6) is maintained by the system (6.6)-(6.7). It is important for the renormalizability of the theory in the PT sense. The difference is that in the QCD the tadpole term remains intact. The essential dynamical source of this difference is that only the tadpole term can generate a mass squared scale parameter. All the other terms present in the skeleton loops decomposition for the full gluon self-energy, seen in Fig. 1, cannot do this 6 , and thus none of their subtracted counterparts with their quark and gluon constants.The system of the relations (6.6)-(6.7) is not only a general but it is a unique one as well. They are not prescriptions, since based on a rigorous tensor algebra derivations, i.e., they are exact mathematical results. The relations (6.6) hold because these constants are related to the corresponding invariant functions for the subtracted counterparts, as described above and shown in the relations (6.5). Evidently, such kind of the relation does not exist for the constant tadpole term $\Delta^{2}_{t}(D)$, that is why it cannot be disregarded at all on this mathematical basis.
All these constants, defined by the corresponding skeleton loop integrals, are QD at the upper limit. They have to be removed/disregarded from the theory on the general basis, i.e., formally put zero, as described above, apart from the tadpole term (2.4). It should remain intact in the QCD full gluon propagator, while in the PT QCD full gluon propagator it has to be removed as well. In what follows all these constants will be called the tadpole-like/type terms, for simplicity. The general question arises now, namely how to understand these exact equalities to zero in the relations (6.6)-(6.7), and hence in the relations (5.5)-(5.6) as well? Each tadpole-like term in these relations is the skeleton loop integral, which thus is a sum of the infinite number of terms. But an equality does not mean that the sum itself is zero. These equalities mean that any tadpole-like term which may appear in the theory by any possible way has to be discarded/disregarded in the theory, i.e, put formally zero, independently from any other tadpole-like terms. In this way such sums will be always zero, indeed. Such constants may appear even not as the result of the subtractions, but, for example as the result of the NL iteration procedure for the full gluon propagator shown in Fig. 1. In this case the tadpole-like terms may be even multiplied by some regularized functions, see appendix A below. So that, whatever their origins would have been in the full gluon propagator, all of them belong to the infinite manifold of the relations (6.6)-(6.7).
Collecting all our results obtained above for the quark and gluon contributions to the gluon self-energy, one obtains
$\Pi^{q}_{\rho\sigma}(q)=T_{\rho\sigma}(q)q^{2}\Pi^{s(q)}_{t}(q^{2}),\quad\Pi^{g}_{\rho\sigma}(q;D)=T_{\rho\sigma}(q)q^{2}\Pi^{s(g)}_{t}(q^{2};D),$ (6.8)
since they become transverse and coincide with their subtracted counterparts. Let us also remind that the invariant functions present in the previous relations are regular functions at small $q^{2}$, i.e., they have no pole-type singularities, and may be only logarithmically divergent at large $q^{2}$. Substituting their sum into the eq. (2.5), one gets
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)$ $+$ $D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)q^{2}\Pi^{s}_{t}(q^{2};D)D_{\sigma\nu}(q)$ (6.9)
$+$ $D^{0}_{\mu\rho}(q)i\delta_{\rho\sigma}\Delta^{2}_{t}(D)D_{\sigma\nu}(q),$
where $\Pi^{s}_{t}(q^{2};D)=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(g)}_{t}(q^{2};D)$.
The corresponding ST identities therefore are not equal to each other, namely $q_{\mu}q_{\nu}(q)D_{\mu\nu}(q)\neq q_{\mu}q_{\nu}D^{0}_{\mu\nu}(q)$, which implies $\xi\neq\xi_{0}$. The gauge-fixing parameter $\xi$ as a function of $\xi_{0}$ is given now by the condition (6.1). Contrary to the gluon SD eq. (5.9), the gluon SD eq. (6.9) is now equivalent to the initial gluon SD eq. (2.5), only being re-written in the different form. That is the reason why we left it denoted as $D$ here and everywhere below. Evidently, the PT gluon SD eq. (5.9) is not equivalent to eq. (2.5). The both full gluon SD eqs. (2.5) and (6.9) by keeping the mass squared parameter explicitly - the tadpole term $\Delta^{2}_{t}(D)$ constitutes the fact that the gauge symmetry of the QCD ground state is not a symmetry of its Lagrangian. The PT renormalizability of the theory seems to be compromised, however see arguments below.
It is important to emphasize once more that we are under an obligation to satisfy at least one of the transverse relations (3.5). Reminding that this is necessary to do in order to get finally to eq. (5.9) in the PT QCD and eq. (6.9) in the QCD. The last one allows to distinguish between the quark $\Delta^{2}_{q}$ and the gluon $\Delta^{2}_{g}(D)$ constants, which are not present in the initial gluon SD eq. (2.5) and the tadpole constant $\Delta^{2}_{t}(D)$, which is explicitly present in it. The satisfied transverse relations have been implemented into the our formalism in a self-consistent way. The fact of the matter is that the system of the relations (6.2)-(6.3) and (6.8)-(6.9) have been uniquely and exactly derived. Then the condition (6.1) can be recovered from the gluon SD eq. (6.9).
7 The QCD full gluon propagator in the generalized gauge
The basic relations to which the full gluon propagator and its free counterpart should satisfy are the corresponding ST identities (3.1) and (3.4), respectively. As emphasized above in section 3, they are important for the general renormalizability of the theory, and thus provide the exact constraints on any solution to QCD 2 . But by themselves they cannot remove the quadratic UV divergences from the theory, as it has been shown above in the relations (6.6)-(6.7), apart from the tadpole term. We have achieved this by formulating the proper subtraction scheme in which the corresponding transverse conditions have been implemented. If some equations, relations or the regularization schemes, etc. do not satisfy the ST identities automatically, i.e., without any additional conditions, then they should be modified and not the identity (3.1). In other words, all the relations, equations, regularization schemes, etc. should be adjusted to it and not vice versa. For example, the transverse condition for the full gluon self-energy (6.2) is not satisfied, but, nevertheless, the general form of the ST identity (3.1) is to be maintained in any case, see eq. (7.3) below. Let us begin with the gluon SD eq. (6.9), which can be equivalently re-written as follows:
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)$ $+$ $D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)\left[q^{2}\Pi(q^{2};D)+\Delta^{2}_{t}(D)\right]D_{\sigma\nu}(q)$ (7.1)
$+$ $D^{0}_{\mu\rho}(q)iL_{\rho\sigma}(q)\Delta^{2}_{t}(D)D_{\sigma\nu}(q),$
saving the tadpole term in the longitudinal component of eq. (6.9), it makes it possible for the tadpole term to appear in the transverse component of the full gluon propagator (7.1) as well. The invariant function $\Pi(q^{2};D)$ is a replacement (for simplicity) of the initial one $\Pi^{s}_{t}(q^{2};D)$ in eq. (6.9). It is worth remanding that it is regular at zero and may have only logarithmic divergences in the PT $q^{2}\rightarrow\infty$ limit. Combining this equation with the decompositions (3.2) and (3.3), one obtains
$d(q^{2})={1\over 1+\Pi(q^{2};D)+(\Delta^{2}_{t}(D)/q^{2})}.$ (7.2)
This relation is not a solution for the full gluon invariant function. It is the NL transcendental equation for the different invariant functions $d(q^{2}),\ \Pi(q^{2};D)$ and the constant $\Delta^{2}_{t}(D)$, i.e., $d=f(D(d))$. Nevertheless, from this expression is clearly seen that in the PT $q^{2}\rightarrow\infty$ regime, the contribution $(\Delta^{2}_{t}(D)/q^{2})$ can be neglected, but the invariant function $\Pi(q^{2};D)$ may still depend on this ratio under the PT logarithms. At the same time, in the NP region of finite and small gluon momenta this term is dominant, and the dependence of $d(q^{2})$ on $\Delta^{2}_{t}(D)$ may be much more complicated due to the transcendental character of eq. (7.2).
Contracting the full gluon SD eq. (7.1) with $q_{\mu}$ and $q_{\nu}$, and substituting its result into the general ST identity (3.1), one arrives at
$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi_{0}\left(1-\xi{\Delta^{2}_{t}(D)\over q^{2}}\right)=i\xi,$ (7.3)
which solution is
$\xi\equiv\xi(q^{2};\xi_{0})={\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)},$ (7.4)
i.e., in this case the gauge-fixing parameter becomes the function $\xi\equiv\xi(q^{2};\xi_{0})$ and thus it is not a constant equal to $\xi_{0}$. Let us point out that the expression (7.4) satisfies condition (6.1), which is as it should be. It is worth emphasizing that the solution (7.4) shows the NL dependence on $\xi_{0}$, which is only possible if gauge symmetry is broken in the ground state. From eq. (7.4) it is clearly seen that in the PT $q^{2}\rightarrow\infty$ regime, which effectively equivalent to the formal $\Delta^{2}_{t}(D)=0$ limit and vice versa, the gauge-fixing parameter $\xi$ goes to $\xi_{0}$, as expected. The generalized gauge (7.4) is the known function of its arguments, while the relation (7.2) is the NL transcendental one. Thus, behind the general inequality $\xi\neq\xi_{0}$ is the constant $\Delta^{2}_{t}(D)$ as its dynamical source, when its contribution can be neglected only then $\xi=\xi_{0}$. The equality holds only for the regularized gluon fields, as underlined above in section 5.
Substituting equations (7.2) and (7.4) into the general decomposition (3.2) for the full gluon propagator, one finally obtains
$D_{\mu\nu}(q)={iT_{\mu\nu}(q)\over q^{2}+q^{2}\Pi(q^{2};D)+\Delta^{2}_{t}(D)}+iL_{\mu\nu}(q){\lambda^{-1}\over q^{2}+\lambda^{-1}\Delta^{2}_{t}(D)},$ (7.5)
[…]
The corresponding ST identity now becomes
$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi(q^{2};\lambda^{-1})=i{\lambda^{-1}q^{2}\over(q^{2}+\lambda^{-1}\Delta^{2}_{t}(D))}.$ (7.6)
This is the generalized gauge since it depends on the constant $\Delta^{2}_{t}(D)$, and when it is zero, one recovers the gauge-fixing parameter for the free gluon propagator. To our best knowledge the generalized gauge (7.6) has been derived for the first time in a such new manner. In the generalized gauge the gauge-fixing parameter $\lambda^{-1}$ is convenient to vary continuously from zero to infinity. The functional dependence of the generalized gauge-fixing parameter $\xi$ (7.4) is fixed up to an arbitrary gauge-fixing parameter $\xi_{0}=\lambda^{-1}$. Unless we fix it, and thus $\xi$ itself, we will call such situation as the general gauge dependence (GGD), see equations (3.1), (7.5), and (7.6). Choosing $\xi_{0}=\lambda^{-1}$ explicitly, we will call such situation as the explicit gauge dependence (EGD). For example, $\xi_{0}=\lambda^{-1}=0$ is called the unitary (Landau) gauge, $\xi_{0}=\lambda^{-1}=1$ is called the t’ Hooft–Feynman gauge, etc. 30 ; 31 ; 32 ; 33 . The formal $\xi_{0}=\lambda^{-1}=\infty$ limit is called as the canonical gauge in 32 . This distinction seems a mere convention, but, nevertheless, it is useful one in QCD because of the presence of a mass scale parameter in its ground state. The generalized gauge requires that there is no other functional expression for $\xi$, apart from given by the relation (7.4) at finite $\xi_{0}=\lambda^{-1}$, in the full gluon propagator (7.5) and the ST identity (7.6) for the regularized gluon fields.
The system of the regularized equations, consisting of the gluon SD eq. (7.1), the expression (7.5) and the ST identity (7.6), constitutes that the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is not a symmetry of its ground state, and it has been exactly and uniquely defined. The tadpole term enters the full gluon self-energy linearly, see the expression (2.2). However, in the full gluon propagator (7.5) it appears in the NL way, because its contribution has been summed up with the help of the gluon SD eq. (7.1). In the PT $q^{2}\rightarrow\infty$ limit, the term $(\Delta^{2}_{t}(D)/q^{2})$ in the expression (7.5) is to be suppressed, and it will be reduced to the PT QCD expression (5.10), reminding that $\xi_{0}=\lambda^{-1}$. The full gluon propagator will behave like the PT QCD gluon propagator at short distance ($q^{2}\rightarrow\infty$), namely $\sim 1/q^{2}$. This is the one of the necessary constraints that a theory is perturbative finite. Other ones such as the corresponding behaviour of the spinor Green’s function, a unitary of $S$-matrix and analyticity (causality) 2 , evidently are beyond the scope of the present paper. Here we have proven that the true dynamical and gauge structures of the QCD/YM ground state are much more complicated than it is required by its Lagrangian’s gauge symmetry. Despite the gauge symmetry is broken in the QCD ground state, its PT renormalizability is not affected within the QCD full gluon propagator in the generalized gauge (7.5), i.e., it has the PT renormalizable behaviour at large $q^{2}$.
8 Discussion
Let us now discuss some interesting features of the explicit presence of the tadpole term in the QCD ground state, i.e., at the fundamental quark-gluon level. First of all, let us underline once more that the tadpole term, existing in the QCD ground state, cannot be discarded in the QCD/YM theory by any means. This was exactly proven in section 6 and used in section 7. Then it will drastically change the structure of the full gluon propagator in the deep IR region $q^{2}\rightarrow 0$. For the sake of transparency and clarity let us use the free gluon propagator in the t’ Hooft-Feynman gauge $\lambda^{-1}=1$, i.e., $D^{0}_{\mu\nu}(q)=i{\delta_{\mu\nu}/q^{2}}$. Then the full gluon propagator (7.1) up to the fist iteration term looks like
$D_{\mu\nu}(q)\sim i{\delta_{\mu\nu}\over q^{2}}-i{T_{\mu\nu}(q)\over q^{2}}\Bigl{[}\Pi(q^{2};D)+{\Delta^{2}_{t}(D)\over q^{2}}\Bigr{]}-iL_{\mu\nu}(q){\Delta^{2}_{t}(D)\over(q^{2})^{2}}+...\ ,$ (8.1)
and it is worth reminding that the invariant function $\Pi(q^{2};D)$ is a regular one of its argument, while all the tadpole-like terms to one loop approximation have been omitted due to the relations (6.6)-(6.7). See appendix A as well, where this iteration is discussed in more detail. Let us also remind that the gluon eq. (7.1) is equivalent to the initial gluon SD eq. (2.1), shown in Fig. 1, and hence to eq. (2.5), only being re-written down in the form suitable for the NP renormalization program to be performed. The higher-order iteration terms will be much more singular with respect to the general ratio $(\Delta^{2}_{t}(D)/q^{2})$, since each iteration invokes additional powers of this ratio. Because of the NL character of the iteration procedure any next iteration will contribute into the previous iteration terms, and thus they become the sums of the infinite number of terms 34 , namely
$D_{\mu\nu}(q)\sim iT_{\mu\nu}{\Delta^{2}_{t}(D)\over(q^{2})^{2}}\sum_{k=0}^{\infty}\Bigl{(}{\Delta^{2}_{t}(D)\over q^{2}}\Bigr{)}^{k}\Phi_{k}(g^{2},\tilde{\lambda}...)+O_{\mu\nu}(1/q^{2}).$ (8.2)
For simplicity, in this iteration expansion is explicitly shown only singular part of the corresponding Laurent series, and thus it represents the summation of all the severe IR singularities (more singular than the PT IR singularity $\sim 1/q^{2}$), which can be accumulated by the full gluon propagator. Evidently, such a summation is only possible due to the existence of the tadpole term. We have shown the singular part, associated with the transverse component of the full gluon propagator, since its longitudinal part is the regular function of the tadpole term, see expression (7.5). The dependence of the coefficients $\Phi_{k}(g^{2},\tilde{\lambda}...)$ on the coupling constant is not known, i.e., such kind of series are the cluster expansions, mentioned above 23 . These coefficients, being the sum of the infinite number of terms by themselves, are resulted from the skeleton loop integrals, contributing into full gluon self-energy through the invariant function $\Pi(q^{2};D)$. Evidently, the iteration ab  infinity leads to the essential singularity at $q^{2}\rightarrow 0$! in the full gluon propagator (8.2), which will remain massless, but depending on the tadpole term. Let us note that the same series can be obtained by expanding the initial NL transcendental relation (7.2) in powers of the ratio $(\Delta^{2}_{t}(D)/q^{2})$.
From the NL iteration expression (8.2) it follows that the full gluon propagator in the IR region ($q^{2}\rightarrow 0$) is dominated by the tadpole contribution, while in the UV region ($q^{2}\rightarrow\infty$) it will be dominated by the free gluon one. This is completely different to QED, where the dressed photon propagator behaves at small photon momentum as its free counterpart ($\sim 1/q^{2}$). This is also true for the full gluon propagator in the PT QCD with unbroken gauge symmetry, investigated in section 5. The principal difference between these theories is due to the existence of the tadpole term in the QCD ground state, which precisely controls the structure of a single full gluon propagator at small $q^{2}$. That is why the dressed PT gluon (5.10) and free gluon states can not appear in the physical spectrum at large distances ($q^{2}\rightarrow 0$), i.e., they will be suppressed in this region by the tadpole term. If this statement will survive the corresponding renormalization program beyond the PT then confinement of the PT and free gluons will be explained. This is important since there is no any other dynamical mechanism to confine these gluon states to the vacuum of QCD or to keep them inside hadrons. To prove the confinement of the singular full gluon propagator itself due to the above-mentioned essential singularity, though possible, but requires completely different investigation within the theory of distributions 26 , and thus is beyond the scope of the present work. Let us note that by the NP program we mean that the full gluon propagator will become free from all the non-physical parameters of the theory in any of its possible solutions (not only in the iteration one).
Eq. (7.5) shows that gluon may acquire mass dynamically if the denominator of this equation becomes zero at some finite point $q^{2}=M^{2}_{g}$. The expression for the free massive vector particle propagator follows in a rather simple way, indeed. For this goal it is necessary to omit the corresponding invariant function and identify the tadpole term with the gluon pole mass $M^{2}_{g}$. In Minkowski metric (by replacing $q^{2}\rightarrow-q^{2}$ and $\delta_{\mu\nu}\rightarrow g_{\mu\nu}$), one gets
$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2}_{g})}\left[g_{\mu\nu}-(1-\lambda^{-1}){q_{\mu}q_{\nu}\over(q^{2}-\lambda^{-1}M^{2}_{g})}\right],$ (8.3)
i.e., it will be expressed in the Stueckelberg gauge 4 . Contracting this expression with respect to $q_{\mu}$ and $q_{\nu}$, one obtains the ST identity (7.6) by replacing there $q^{2}\rightarrow-q^{2}$. In other words, the Stueckelberg gauge (8.3) is a particular case of the generalized gauge (7.6). The PT QCD full gluon propagator (5.10) cannot generate the gluon pole mass at some finite point, since it remains always massless. The full gluon propagator (7.5) provides a rather convenient framework in order to formulate and further develop the NP renormalization program for the massive gluon fields. This program will suppose to show that the full gluon propagator (7.5), and hence its free massive counterpart (8.3), can not be put on the mass-shell $q^{2}=M^{2}_{g}$. Acquiring mass dynamically, the gluon, nevertheless, will remain the off-mass-shell object. So that it will not appear in the physical spectrum (confinement of the massive gluon fields within our approach).
In the presence of the mass squared scale parameter - the tadpole term - the role of the QCD coupling constant $g^{2}$ becomes unimportant. This is also evidence of the ’dimensional transmutation’, $g^{2}\rightarrow\Delta^{2}_{t}(D)$ 2 ; 35 ; 36 , which occurs whenever a massless theory acquires mass dynamically. It is a general feature of spontaneously symmetry breaking in field theories. We distinguish between the PT and full QCD by the explicit presence of the tadpole term in the latter one, and not by the magnitude of the coupling constant even at the regularized gluon fields level yet. In the both cases the gluon fields remain strongly interacted, apart from the asymptotic freedom (AF) regime 37 ; 38 .
Any deviation of the full gluon propagator from the free one in QCD requires the presence of a mass squared scale parameter. Even in the AF regime there is a scale breaking,
$D_{\mu\nu}(q)\sim\delta_{\mu\nu}\Bigl{[}{ig^{2}\over 1+g^{2}b_{0}\ln(q^{2}/\Lambda^{2}_{QCD})}\Bigr{]}(1/q^{2}),\quad q^{2}\rightarrow\infty,$ (8.4)
which has been discovered many years ago 37 ; 38 . When $q^{2}\rightarrow\infty$ the dependence on the coupling constant $g^{2}$ vanishes, indeed. The expression (8.4) is written down in the t’ Hooft-Feynman gauge and in the approximation of the main PT logarithms with the colour group factor $b_{0}>0$. $\Lambda^{2}_{QCD}$ determines the scale of the non-trivial PT dynamics in the QCD. However, in section 5, we have already shown that the PT cannot generate the scale parameter, having the dimension of a mass squared. From the renormalization group equations analysis, it follows that any mass to which can be assigned any physical meaning is exponentially suppressed in the weak coupling limit to every order of the PT 2 ; 35 , namely
$M\sim\mu\exp(-1/b_{0}g^{2}),\quad\quad g^{2}\rightarrow 0,$ (8.5)
where $\mu$ is the arbitrary renormalization point. None a finite mass can survive in the PT weak coupling limit or, equivalently, in the PT $q^{2}\rightarrow\infty$ regime, indeed. So the question where the finite mass comes from? cannot be answered by the PT! It is entirely of the NP origin! It has to come from the IR region which is controlled by the tadpole term, as discusses just above in this section. Its finite version should be somehow related to $\Lambda^{2}_{QCD}$ (reminding that the ratio $(\Delta^{2}_{t}(D)/q^{2})$ may appear under the PT logarithms).
9 Conclusions
Our formalism makes it possible to remove from the theory the QD constants of the gluon degrees of freedom in the same way as the QD constants of the quark degrees of freedom. This has lead to the discovery that in addition to the PT gluon field configurations in the QCD ground state (more precisely in its YM sector), preserving the exact gauge symmetry of its Lagrangian, there also exists such a type of the gluon field configurations which do not maintain this symmetry. But the PT renormalizable properties of the theory will not be affected. The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared. It is dynamically generated by the self-interaction of the multiple massless gluon modes involving only the point-like 4-gluon vertex, see eq. (2.4) and Fig. 1. The full 4-gluon vertex survives when all the gluon momenta involved go to zero, becoming thus its point-like counterpart, while the full 3-gluon vertex does not survive in this limit 10 ; 23 ; 39 . By the design the tadpole term is present in the QCD ground state from the very beginning, i.e., not introducing by hand. It is only one which explicitly contributes into the skeleton loops decomposition of the full gluon propagator (shown in Fig. 1), while all the other tadpole-like terms don’t. Unlike them, being the QD but already regularized constant, the tadpole term is not connected to any external gluon momentum, and thus may generate a mass squared scale parameter 6 , indeed. Thus, there is no need to extract such a scale parameter in the QCD vacuum in a some other way. In 10 the importance of the tadpole term has been recognized and exploited on an intuitive basis rather than a logical one. However, from now on its existence cannot be undermined. Besides its combinatorial (mathematical) meaning it has a deep dynamical one as a source of the gauge symmetry violation in the QCD vacuum. Its renormalized (i.e., finite) version will determine a scale of the NP dynamics in QCD and in its true ground state.
For the both types of the gluon field configurations have been derived the corresponding expressions for the full gluon propagators describing their propagation in the vacuum of QCD. For the second type the corresponding NP QCD full gluon propagator (7.5) is written down in the generalized gauge (7.6). From these expressions it is clearly seen that in the PT $q^{2}\rightarrow\infty$ limit eq. (7.5) coincides with eq. (5.10). This is an evidence that the PT renormalizable properties of the theory will not be affected by the new type of the gluon field configurations. In other words, the renormalizability of the theory will be maintained within our approach to the full gluon propagator. It is worth emphasizing once more that the massive and severely singular gluon field configurations can be described by eq. (7.5) only, while eq. (5.10) describes the propagation of the massless gluon modes with the PT IR singularities $\sim 1/q^{2}$ only. Our ground-breaking findings are as follows:
a. One of the transverse relations (3.5) is to be necessarily satisfied, in order to make the theory renormalizable in the PT sense.
b. The subtraction scheme has been formulated, which makes it possible to remove from theory all the quadratically UV divergent constants on the general basis.
c. The exact gauge symmetry of the YM Lagrangian is not a symmetry of its ground state. The dynamical source of this effect is the tadpole term. If it formally put zero then these gauge symmetries coincide.
d. The system of the relations (6.6)-(6.7) are exact mathematical results, i.e., they are not prescriptions.
e. The new NP approach to the YM theory and its ground state has been formulated, by keeping the corresponding ST identities always valid. It is based on the system of eqs.  (7.5)-(7.6), which is defined up to the corresponding invariant function with the known asymptotic properties. Eqs. (7.5)-(7.6) are suitable for the NP renormalization program to be performed.
f. The NL dependence $\xi$ on $\xi_{0}$ in the generalized gauge (7.4) in the PT $q^{2}\rightarrow\infty$ or, equivalently, formal $\Delta^{2}_{t}(D)=0$ limits becomes linear one $\xi=\xi_{0}$.
g. The DRM 6 ; 25 ; 26 ; 27 prescription to simply ignore all the QD constants (including the tadpole term) in the PT QCD has been put on a firm mathematical ground from now on, see eqs. (5.4)-(5.6).
In obtaining all these results no specific regularization schemes have been used. Neither special gauge choice nor any truncations, approximations and assumptions have been made either in dealing with the regularized skeleton loop integrals, contributing to the gluon self-energy. The exact analytic derivations have been done, based on a rigorous tensor algebra rules, such as the decomposition into the independent structures, the subtractions, etc.
10 Summary
It is shown in the most general and unique way that the gauge symmetry of the QCD Lagrangian is broken in its ground state, indeed. This has been achieved without solving the full gluon SD eq. (2.5), but recovering only its true dynamical and gauge structures in more detail in the framework of our formalism. So that, the significant peace of the information on these important issues was not missed within our general approach. Its breakthrough results can be summarized in terms of the following theorem:
Tadpole term existence and the true QCD ground state: Prove that in the QCD ground state exist such gluon field configurations which break/violet the gauge symmetry of its Lagrangian, but the PT renormalizability of the theory is maintained. Existence of this effect includes establishing the true role of the tadpole term in the QCD ground state.
Just this theorem has been proven above in this paper, which resolves the dilemma formulated at the end of section 2. In the proof of this theorem we have shown how to save the tadpole term in QCD and its ground state, without affecting the PT renormalizable properties of the theory, while removing all the other QD constants from it. The new derived solution with the gauge symmetry breaking will apparently satisfy the bounder conditions in QCD at the fundamental quark-gluon level: confinement at large distances ($q^{2}\rightarrow 0$) and AF at short distances ($q^{2}\rightarrow\infty$), as discussed in section 8. The solution with the gauge symmetry preservation can not satisfy them. All this underlines once more the importance of the theorem above.
The perspectives for further work are as follows: to formulate the NP multiplicative (MP) renormalization programs for the full gluon propagator (7.5), which is only regularized one at this stage. Unfortunately, this is not a simple task, since such kind of a programs have to be different for the massive and severely singular gluon field configurations, briefly discussed in section 7. For this purpose it is necessary to take into account the transcendental and NL structure of eq. (7.2). They are also needed to be formulated in the most general and unique ways. To ensure a unitary of $S$-matrix elements by these NP MP renormalization programs should be requested as well.
To understand a mass dynamical generation at the quark-gluon level is the first necessary step to understand the existence of the physical mass spectrum at the hadronic level in QCD as a theory of strong interactions. The confinement first and only after the PCAC phase transitions in QCD have to be understood 2 ; 23 . QCD is a self-consistent quantum gauge field theory. It does not need any extra degrees of freedom (such as Higgs fields 40 ) in order to generate a mass at the fundamental quark-gluon level.
Concluding, a few remarks are in order. The wide-spread opinion that the gauge symmetry cannot be broken/violated is misleading. The Elitzur’s theorem 41 has been proven in lattice field theory. It is supposed to be valid for the continuous theories as well, but if yes than only for the abelian systems, like QED. To the non-abelian systems, such as the YM fields, this theorem cannot be applied. In this paper we have just proven that the gauge symmetry of the YM theory is broken in its vacuum. The relevant existing knowledge on a mass dynamical generation and spontaneously symmetry breaking subjects has been taken into account in the list of publications, presented in our paper (and see references in each of the cited items there).
Acknowledgments
The authors are grateful to P. Levai, P. Forgács, J. Nyiri, T.S. Biró, Gy. Kluge, M. Vasúth for useful suggestions, remarks, discussions and help. The work was supported by the Hungarian National Research, Development and Innovation Office (NKFIH) under the contract numbers OTKA K135515 and NKFIH 2019-2.1.11-TÉT-2019-00078, 2019-2.1.11-TÉT-2019-00050, the Wigner Scientific Computing Laboratory.
Appendix A The full gluon propagator up to one loop
Figure 2: The SD equation for the full gluon propagator up to one loop as present in 6 . It is useful to compare it with Fig. 1 and its description in the text after it.
In order to explicitly compare the tadpole term (2.4) with the other tadpole-type terms let us consider the iteration of the gluon SD eq. (2.1), shown in Fig. 1, up to one loop, shown in Fig. A.2. The last two terms in this figure appear because of the NL nature of the iteration series, when the next iterations contribute into the previous ones. That is the reason why such multiplied tadpole-type terms are not present in the skeleton loop decomposition for the full gluon propagator in Fig. 1, but have to be shown up in the iteration terms, seen in Fig. A.2. This iteration has been exactly calculated in 6 by using the DRM, while rightfully now ignoring the tadpole term itself and all the other tadpole-likes ones, in accordance with the relations (5.5)-(5.6). It is worth emphasizing that in 6 this iteration has not been called the PT expansion up to $O(g^{2})$, though looks like it, but expansion up to $O(\hbar)$. It is well-known that in QCD the expansion in powers of the coupling constant does not make any sense because it is strong. The PT expansion in QCD makes sense only in the AF 37 ; 38 regime when it becomes weak. In order to perform analytical derivations below, we will take the interaction vertices as the point-like ones. The external gluon propagators will be taken as the free ones. All the skeleton loop propagators will remain as the full ones. Then the corresponding iteration will look like Fig. A.2, and hence can be called as quasi-one loop expansion, discussed in some details in section 8, when it was called as the cluster one. Note that below we do not show explicitly the dependence on the coupling constant, since our skeleton loop quark, gluon and ghost propagators are the full ones, so that the exact dependence on the coupling constant of the expressions below is not known. Also we will omit below all the colour group factors and other overall numerical numbers, since all of them are unimportant here.
However, it is instructive first to show explicitly how works our subtraction scheme, using only the third diagram in the second line of Fig. A.2, for simplicity. From 6 by replacement $\mu\rightarrow\rho$ and $\nu\rightarrow\sigma$ and fixing $\omega=2$ and $\delta_{\rho\rho}=4$ in eq. (6.22) there, one finally obtains an analytical expression for this term as follows:
$\Pi^{(3)}_{\rho\sigma}(q)\sim\int d^{4}l\ {N_{\rho\sigma}(l,q)\over l^{2}(l+q)^{2}},$ (A.1)
where, apart from below, we put the gluon loop propagators as free ones in the t’ Hooft-Feynman gauge as well, for the sake of transparency of our derivations for this diagram. The tensor function $N_{\rho\sigma}(l,q)$ is
$N_{\rho\sigma}(l,q)=10l_{\rho}l_{\sigma}+5(l_{\rho}q_{\sigma}+q_{\rho}l_{\sigma})-2q_{\rho}q_{\sigma}+(5q^{2}+2(lq)+2l^{2})\delta_{\rho\sigma}.$ (A.2)
Putting in eqs. (A.1) and (A.2) $q=0$, one gets
$\Pi^{(3)}_{\rho\sigma}(0)\sim\int d^{4}l\ {N_{\rho\sigma}(l,0)\over(l^{2})^{2}}=\int d^{4}l\ {(10l_{\rho}l_{\sigma}+2l^{2}\delta_{\rho\sigma})\over(l^{2})^{2}},$ (A.3)
and, obviously, this integral is quadratically divergent at the upper limit, but is assumed to be regularized from above and below from the very beginning, as emphasized throughout this work.
Let us now compose the subtracted counterpart of the initial integral (A.1), in accordance with the subtractions (4.1), as follows:
$\Pi^{(3s)}_{\rho\sigma}(q)=\Pi^{(3)}_{\rho\sigma}(q)-\Pi^{(3)}_{\rho\sigma}(0)\sim\int d^{4}l\ {N^{s}_{\rho\sigma}(l,q)\over(l^{2})^{2}(l+q)^{2}},$ (A.4)
where
$N^{s}_{\rho\sigma}(l,q)$ $=$ $5l^{2}(l_{\rho}q_{\sigma}+q_{\rho}l_{\sigma})-20(ql)l_{\rho}l_{\sigma}+l^{2}(q^{2}-2(lq))\delta_{\rho\sigma}$ (A.5)
-- $2l^{2}q_{\rho}q_{\sigma}-10q^{2}l_{\rho}l_{\sigma}.$
Because of the two last terms and term $l^{2}q^{2}\delta_{\rho\sigma}$ in this equation, the subtracted integral (A.4) is only logarithmically divergent at the upper limit. So that the integral (A.3) is one of the tadpole-like terms which should be disregarded on the general basis due to the relations (6.6)-(6.7) within our approach. The subtracted integral (A.4) finally will contribute to the invariant function which appear in the full gluon propagator (7.5).
The tadpole term up to one skeleton loop, shown first in the third line of Fig. A.2, analytically can be written down as follows:
$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lT^{0}_{\rho\sigma\alpha\beta}D_{\alpha\beta}(l)=\delta_{\rho\sigma}\Delta^{2}_{t}(D).$ (A.6)
Being the QD but already regularized constant, it is not connected to any external gluon momentum. In this respect it is different from all other tadpole-like terms which appear as a result of the subtraction scheme or in the NL iteration procedure. Also it directly generate the mass squared parameter, and thus explicitly violets the gauge symmetry of the QCD Lagrangian in its ground state 6 . We have shown that in the PT QCD it should be neglected on the general basis as well as all other tadpole-type terms. But in the QCD it should remain intact, while all other tadpole-types terms have to be removed from the theory, as it has been proven in section 6.
Let us now investigate the tadpole-like term with the gluon skeleton loop, shown second in the third line of Fig. A.2. Its analytical expression is
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l\ T^{0}_{\nu\zeta\mu}(l,p)D_{\nu\zeta}(l),$ (A.7)
where the gluon momentum $p$ is, in fact, zero, i.e., $p=0$ but it is convenient to go to this limit at the final step only. Both gluon propagators in the t’ Hooft-Feynman gauge are $D^{0}_{\mu\mu^{\prime}}(p)=\delta_{\mu\mu^{\prime}}/p^{2}$ and $D_{\nu\zeta}(l)=\delta_{\nu\zeta}d(l^{2})/l^{2}$, where $d(l^{2})$ is the corresponding invariant function of the full gluon propagator. Let us remind that such kind of the skeleton loop integrals are assumed to be regularized from above and below. Using now the Euclidean space Feynman rules for the corresponding vertices present in 6 , one arrives at
$T^{0}_{\rho\sigma\mu^{\prime}}(q,p)$ $=$ $-(2q+p)_{\mu^{\prime}}\delta_{\rho\sigma}+(q-p)_{\sigma}\delta_{\rho\mu^{\prime}}+(2p+q)_{\rho}\delta_{\sigma\mu^{\prime}},$
$T^{0}_{\nu\zeta\mu}(l,p)$ $=$ $2(l+p)_{\mu}\delta_{\nu\zeta}-l_{\zeta}\delta_{\nu\mu}-(2p+l)_{\nu}\delta_{\zeta\mu}.$ (A.8)
Substituting these expressions into the eq. (A.7), one finally obtains
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p){2\over p^{2}}\int d^{4}l{d(l^{2})\over l^{2}}[3l_{\mu}+p_{\mu}].$ (A.9)
By the symmetry integration ($d^{4}l=l^{3}dl=(1/2)l^{2}dl^{2}$ and all the overall numerical numbers due to the integration over angular variables in four-dimensional Euclidean space 6 omitted) the first loop integral is zero, while the second one leads to
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p)p_{\mu}{2\over p^{2}}\times\tilde{\Delta}^{2}_{g}(d),$ (A.10)
where $\tilde{\Delta}^{2}_{g}(d)=\int d^{4}l(d(l^{2})/l^{2})$ is the QD but regularized skeleton loop integral, having the dimension of a mass squared. It is one of the constants present in the relations (6.7)-(6.8). It should be discarded on the general basis within our approach, i.e., put formally zero, so that $\tilde{\Delta}^{2}_{g}(d)=0$, even before the final $p^{2}=-M^{2}\rightarrow 0$ limit. Let us note that the cancellation of the pole $1/p^{2}=-1/M^{2}$ in eq. (B.10) can be always achieved by going to the dimensionless variable $x=l/M$ in the skeleton loop integral (A.9), i.e., this pole is not a problem here. Thus, $\Pi^{(4)}_{\rho\sigma}(q,p)=0$.
Let us finally investigate the tadpole-like term with the ghost skeleton loop, shown third in the third line of Fig. A.2. In analogy with eq. (A.7) its analytical expression is
$\Pi^{(5)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l{l_{\mu}\over l^{2}}d_{gh}(l^{2}),$ (A.11)
where $d_{gh}(l^{2})$ is the invariant function for the full ghost propagator $G(l)$. The vertex $T^{0}_{\rho\sigma\mu^{\prime}}(q,p)$ is explicitly shown in the relations (A.8), and the remaining free gluon propagator is also the same. The altered rule 6 for the gluon-ghost vertex gives $(1/2)[(l_{\mu}+l{\mu})=l_{\mu}$ and the gluon momentum $p$ is zero again at the final stage. Since the nominator of this skeleton loop integral contains the loop variable linearly, then by the symmetry integration it is simply zero from the very beginning, i.e., $\Pi^{(5)}_{\rho\sigma}(q,p)=0$.
Concluding, neither the tadpole-like term with the gluon skeleton loop nor its ghost counterpart contribute into the full gluon propagator. In these cases the subtractions with respect to the external gluon momentum $q$ (though possible), but were not even necessary to perform, since the final results would have been the same zeros. The primary goal of this work is to prove that the tadpole term cannot be disregarded by any means. All the other tadpole-like terms are to be always discarded on the general basis, i.e., put zeros, as explicitly demonstrated here as well. This was not a prescription, but being an exact mathematical result within our formalism.