### v3

# The gauge symmetry violation in the QCD ground state

Abstract
The purpose of this work is to report on convincing progress made in the investigation of the true dynamical and gauge structures of the QCD ground state (vacuum). It also sheds new significant insights into a mass squared dynamical generation in it. For this purpose, we use only the tensor algebra derivation rigorous rules in dealing with the corresponding skeleton loop integrals, contributing into the full gluon self-energy. The novel constraint on any solution to QCD has been exactly and uniquely derived in the framework of the the Slavnov-Taylor identities for the gauge particles propagators. Its general solution makes it possible to prove that the gauge symmetry of the QCD Lagrangian is broken in its ground state. The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared. The exact proof has been given that it cannot be disregarded from the theory and its ground state by any means. The gauge symmetry of the QCD Lagrangian forbids such kind of terms, but the perturbative renormalizability of the theory will not be affected. The particular solution of the novel constraint describes the gluon field configurations with unbroken gauge symmetry in the QCD vacuum. The solution with broken gauge symmetry dominates over the solution with unbroken one at large distances, while being in agreement with it at short distances.
keywords:
Theoretical particle physics, Quantum Chromodynamics, ground state, full gluon self-energy, gauge
PACS:
11.10.-z , 11.15.-q , 12.38.-t , 12.38.Lg , 12.38.Aw
[…]
[table]capposition=top
1 Introduction
The quark model (QM) treats the strongly-interacting hadrons (baryons and mesons) as bound-states of quarks, by emitting and absorbing gluons. The theory, which purpose is to describe the properties of the observed hadrons in terms of the non-observable quarks and gluons from first principles, is Quantum Chromodynamics (QCD) [1, 2, 3, 4, 5, 6, 7, 8, 9]. It is widely accepted as the quantum gauge field theory of strong interactions. Being such a theory, it has to undergo the two phase transitions: the first one at the fundamental (microscopic) quark-gluon level is – the confinement phase transition – in order to explain why all the physical states are colour-singlets. The second one at the (macroscopic) hadronic level is – the PCAC (Partially Conserved Axial Currents) phase transition – in order to explain the soft pion physics [2, 23].
However, this purpose remains a formidable task yet because of the multiple dynamical and topological complexities of low-energy particles (hadrons) physics, originated from QCD and its ground state (vacuum). This happens because QCD as a gauge theory of the gluon-gluon and quark-gluon strong interactions suffers from a few important conceptual problems. Here we focus on the two of them, being in the close connection with each other.
A. Whether the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is a symmetry of its ground state or not?
B. The dynamical generation of a mass squared in the vacuum of QCD, since the gauge invariance of its Lagrangian forbids such kind of terms, for example like the gluon mass term $M^{2}_{g}A_{\mu}A_{\mu}$.
The solutions of these fundamental problems is a key to the explanation of the first confinement phase transition in QCD. The properties and symmetries of the QCD Lagrangian, and thus including its Yang – Mills (YM) part, are well-known [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The propagation of gluons is one of the main dynamical effects in the QCD ground state. The importance of the corresponding equation of motion is due to the fact that its investigation is supposed to reflect the dynamical and gauge structures of the QCD ground state. The gluon Schwinger – Dyson (SD) equation [2, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] is highly non-linear (NL) because of the self-interaction of massless gluon modes. They are taken into account by the full gluon self-energy, which is the main part of the gluon SD equation itself. We will show that the general properties of the full gluon self-energy point out on some new dynamical and gauge aspects of the true QCD ground state. Due to these impacts, the novel insights into a mass dynamical generation in the QCD vacuum are also present within proposed below formalism. It is based on the tensor algebra derivation rigorous rules only, which is widely using in the theoretical and mathematical physics.
2 Gluon SD equation
The system of the SD equations of motions in QCD, including the above-mentioned gluon SD equation, in general, contains much more informations on the properties of the theory than its Lagrangian can provide at all [2]. So it might be the place where a mas squared scale parameter, forbidden by the QCD Lagrangian gauge symmetry, could be generated. Among them, the special role belongs to the equations, describing the propagation of the fundamental constituents in the QCD vacuum – gluons and quarks. The gluon SD equation has a rather complicated tensor structure because of its above-mentioned NL character. Let us emphasize that all the known interactions in nature (electroweak, gravitational and strong) are described by the gauge theories. The common mathematical language for them is the tensor algebra, which is only one used throughout this paper. That is why our works will be easily understood by a broad spectrum of researches from the different areas of the theoretical physics. The gluon SD equation analytically looks like
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{0}_{\mu\rho}(q)i\Pi_{\rho\sigma}(q;D)D_{\sigma\nu}(q),$ (2.1)
where $D_{\mu\nu}(q)$ and $D^{0}_{\mu\nu}(q)$ denote the full gluon propagator and its free counterpart, respectively. $\Pi_{\rho\sigma}(q;D)$ is the full gluon self-energy which depends on the full gluon propagator due to the non-abeian character of QCD. Here and everywhere below we omit the colour group indices, for simplicity, because of their final factorization, for example $D^{ab}_{\mu\nu}(q)=D_{\mu\nu}(q)\delta^{ab}$. Eq. (2.1) in terms of the corresponding skeleton loop diagrams is shown in Fig. 1.
Figure 1: The SDE for the full gluon propagator, as present in [10].
Here stringy lines are for the free gluon propagator, while $D$ denotes its full counterpart. $S$ with solid lines denotes the full quark propagator, and $\Gamma$ denotes the full quark-gluon vertex. $G$ with dashed lines denotes the full ghost propagator, and $G_{\mu}$ is the full ghost-gluon vertex. Finally, $T_{3}$ and $T_{4}$ denote the full 3- and 4-gluon vertices, respectively.
The full gluon self-energy is convenient to present as the sum of the three independent terms, namely
$\Pi_{\rho\sigma}(q;D)=\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)+\Pi_{\rho\sigma}^{t}(D),$ (2.2)
where $\Pi^{q}_{\rho\sigma}(q)$ describes the skeleton loop contribution of the quark degrees of freedom as an analogue to the vacuum polarization tensor in Quantum Electrodynamics (QED) [7, 21]. Here and below the superscript or subscript ’$q$’ means quark (not to be mixed up with the gluon momentum variable $q$). The gluon part of the full gluon self-energy by itself is the sum of a few independent terms as follows:
$\Pi^{g}_{\rho\sigma}(q;D)$ $=$ $\Pi^{gh}_{\rho\sigma}(q)+\Pi^{(1)}_{\rho\sigma}(q;D^{2})$ (2.3)
$+$ $\Pi^{(2)}_{\rho\sigma}(q;D^{4})+\Pi^{(2^{\prime})}_{\rho\sigma}(q;D^{3}),$
and $\Pi^{gh}_{\rho\sigma}(q)$ describes the skeleton loop contribution associated with the ghost degrees of freedom. $\Pi^{(1)}_{\rho\sigma}(q;D^{2})$ represents the skeleton loop contribution, containing the 3-gluon vertices only. Finally, $\Pi^{(2)}_{\rho\sigma}(q;D^{4})$ and $\Pi^{(2^{\prime})}_{\rho\sigma}(q;D^{3})$ describe the skeleton two-loop contributions, which combine the 3- and 4-gluon vertices. All these quantities are given by the corresponding skeleton loop diagrams in Fig. 1, and they are independent from each other. The last four terms explicitly contain the full gluon propagators in the corresponding powers symbolically shown above. The analytical expressions for the corresponding skeleton loop integrals [22], in which the symmetry and combinatorial coefficients and signs have been included, are not important here. We are not going to calculate any of them explicitly, and thus to introduce into them any truncations/approximations/assumptions or choose some special gauge. Any skeleton loop integral in Fig. 1 is the sum of the infinite number of terms. Moreover, the full vertices entering these skeleton loop integrals are themselves determined by an infinite series of the corresponding multi-loop skeleton terms [2, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. In these NL series the dependence on the coupling constant may not be simple, i.e., in fact, it is not known. Such kind of series are the so-called ’cluster’ expansions [23], indeed.
The constant tadpole term $\Pi_{\rho\sigma}^{t}(D)$ is defined as,
$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lD_{\alpha\beta}(l)T^{0}_{\rho\sigma\alpha\beta}=g_{\rho\sigma}\Delta^{2}_{t}(D),$ (2.4)
where $g_{\rho\sigma}=T_{\rho\sigma}(q)+L_{\rho\sigma}(q)=T_{\rho\sigma}(q)+q_{\rho}q_{\sigma}/q^{2}$. In what follows the constant $\Delta^{2}_{t}(D)$ will be called as the tadpole term itself, since it deviates from the definition (2.4) by the tensor $g_{\rho\sigma}$ only.
It is instructive for further purpose to present the initial gluon SD eq. (2.1), on account of the expressions (2.2) and (2.4) as follows:
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{T}_{\mu\rho}(q)+D^{L}_{\mu\rho}(q)$ (2.5)
with
$D^{T}_{\mu\nu}(q)$ $=$ $D^{0}_{\mu\rho}(q)i\left[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)\right]D_{\sigma\nu}(q),$
$D^{L}_{\mu\nu}(q)$ $=$ $D^{0}_{\mu\rho}(q)ig_{\rho\sigma}\Delta^{2}_{t}(D)D_{\sigma\nu}(q).$ (2.6)
All the terms which contribute to the full gluon self-energy eq. (2.2), and hence eq. (2.5), are tensors, having the dimensions of a mass squared. All these skeleton loop integrals are therefore quadratically divergent (QD), i.e., ultraviolet (UV) divergent in the perturbation theory (PT) regime, and so they are assumed to be regularized from above. Contrary to QED, QCD being a non-abelian gauge theory can suffer from the severe infrared (IR) singularities in the $q^{2}\rightarrow 0$ limit (i.e., more singular than the PT one, which is proportional to $\sim 1/q^{2}$). Evidently, they will be generated due to the self-interaction of massless gluon modes. Thus, all the skeleton loop integrals, contributing to the full gluon self-energy, are assumed to be regularized from below as well. This means that all the expressions are regularized (i.e., we render infinite skeleton loop integrals to be finite). How precisely the regulating parameters have to be introduced is not important here, since we are not going to calculate any of these skeleton loop integrals explicitly, as emphasized above. They will be assumed but not shown in all regularized expressions, for simplicity.
The tadpole term $\Delta^{2}_{t}(D)$, defined in eq. (2.4), is nothing else but the QD skeleton loop integral, which is already regularized from below and above as well as all the other such kind of constants which may appear in the theory by any possible ways (the subtraction schemes, the iteration solutions, etc., see below). Being a mass squared regularized quantity, it is explicitly present in the QCD ground state, see eq. (2.5). However, it violets the gauge symmetry of the QCD Lagrangian. Then a natural question arises why is it present in the vacuum of QCD at all if it makes the theory to look non-renormalizable from the very beginning? The standard solution to this dilemma was to remove this and all other QD constants in order to make theory renormalizable in accordance with the QCD Lagrangian’s exact gauge symmetry. However, here we are going to solve this dilemma by asking the question is it possible to retain the tadpole term in the QCD vacuum, but without affecting the PT renormalization properties of the theory? This is important to ask because, unlike to other QD constants, the tadpole term is only one which dynamically generates a mass squared scale parameter, dominated by the point-like four-gluon vertex only, and thus does not depending on the external gluon momentum [6]. Also, one of our primary aims will be to clarify its role in the dynamical and gauge structures of the QCD vacuum in more detail.
Concluding, let us remind that the whole gluon momentum range is $q^{2}\in[0,\infty)$. In what follows we will work in Euclidean metric $q^{2}=q^{2}_{0}+\bf q^{2}$ since it implies $q_{i}\rightarrow 0$ when $q^{2}\rightarrow 0$ and vice versa. This makes it possible to avoid the unphysical IR singularities on the light cone.
3 Exact constraint on any solution to QCD
It is well known that all the Slavnov-Taylor (ST) identities which appear in QCD [1] are important for the renormalizability of the theory, and thus they ”are exact constraints on any solution to QCD” [2]. Among them the ST identities [2, 24, 35] for the gauge particle propagators play a dominant role, since they determine the gauge structure of the QCD ground state (vacuum).
The ST identity for the full gluon propagator looks like
$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi,$ (3.7)
where $\xi$ is the gauge-fixing parameter. It implies that the general tensor decomposition of the full gluon propagator in the covariant gauge is as follows:
$D_{\mu\nu}(q)=i\left[T_{\mu\nu}(q)d(q^{2})+\xi L_{\mu\nu}(q)\right]{1\over q^{2}},$ (3.8)
where the invariant function $d(q^{2})$ is the corresponding Lorentz structure of the full gluon propagator (or the gluon invariant function). Throughout this paper we use the standard definition of $T_{\mu\nu}(q)=\delta_{\mu\nu}-q_{\mu}q_{\nu}/q^{2}=\delta_{\mu\nu}-L_{\mu\nu}(q)$ in Euclidean metric. Any invariant functions associated with the projection operators $T_{\mu\nu}(q)$ and $\delta_{\mu\nu}$ are the same, and thus $D_{\mu\nu}(q)$ is defined up to its longitudinal part $L_{\mu\nu}(q)$. This should be also valid for its equation of motion. By formally setting $d(q^{2})=1$ in eq. (3.8), then one obtains the free gluon propagator
$D^{0}_{\mu\nu}(q)=i\left[T_{\mu\nu}(q)+\xi_{0}L_{\mu\nu}(q)\right](1/q^{2}),$ (3.9)
where $\xi_{0}$ is the corresponding gauge-fixing parameter. The general ST identity (3.7) will look like
$q_{\mu}q_{\nu}D^{0}_{\mu\nu}(q)=i\xi_{0}.$ (3.10)
Let us note in advance that from the very beginning the gauge-fixing parameter for the full gluon propagator $\xi$ is not to be identified with its free counterpart $\xi_{0}$. Since the gauge freedom in QCD is determined by $\xi_{0}$, there should exist a relation between them.
Let us now perform some exact algebraic derivations which are necessary for the further purposes. Contracting the full gluon SD eq. (2.1) with $q_{\mu}$ and $q_{\nu}$, on account of the relations (3.7)-(3.10), one gets
$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2},$ (3.11)
where the full gluon self-energy, due to the eqs. (2.2) and (2.4), can be present as follows:
$\Pi_{\rho\sigma}(q;D)=\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)+\delta_{\rho\sigma}\Delta^{2}_{t}(D),$ (3.12)
and contracting it with with $q_{\rho}$ and $q_{\sigma}$, and because of the previous relation (3.11), one obtains
$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}-q^{2}\Delta^{2}_{t}(D).$ (3.13)
Obviously, this transverse relation can be obtained by the corresponding contraction of eq. (2.5). It explicitly depends on the tadpole term $\Delta^{2}_{t}(D)$. If it is formally put zero, i.e., by hand, then the both transverse relations (3.11) and (3.13) coincide with each other. This underlines its important role in the QCD ground state, which one precisely is to be clarified in our paper. The transverse relation (3.11) and the transverse relation obtained from eq. (3.12) are independent from each other, since the first one has been derived by using the gluon SD eq. (2.2), while the second one by using the initial definition (2.6). From all these transverse relations one concludes that by themselves they cannot remove the QD constants which may appear in the theory (in order to make it in the PT sense renormalizable), and to fix the relation between $\xi$ and $\xi_{0}$.
Let us now remind that in QCD the quark skeleton loop contribution into the full gluon self-energy can be always made transverse because of the electric charge conservation which flows around the closed loop (in complete analogy with QED, see [10] and derivations below). Therefore, $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}=0$ and this relation holds independently from the gluon contribution, which has been defined in eq. (2.3). At the same time, it is well-known that in QCD that just the Faddeev – Popov (FP) ghost contribution [43], $\Pi^{gh}_{\rho\sigma}(q)$ makes the transverse relation for this term satisfied (independently from the tadpole term (2.4), which itself has the transverse projection operator structure). It is instructive to emphasize that each term in the sum (2.3) cannot be made transverse, only the whole sum. How explicitly this works can be found in any text book on QCD. For the most transparent calculations up to one loop contribution to the transverse gluon self-energy (2.3) see for example [3, 4, 5, 6]. Of course, such a transversality will be achieved in the higher loop iteration terms as well, i.,e, it is a general fact, as it is requested by the satisfied transverse relation for the (2.3) term. It is a sum of the skeleton loop contributions, as underlined from the very beginning in this work. This means that $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$ and this relation holds independently from the quark contribution.
So that, in general, in QCD one has
$q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0,$ (3.14)
and then from the relation (3.13), one immediately arrives at the exact condition, namely
${(\xi_{0}-\xi)\over\xi\xi_{0}}q^{2}=\Delta^{2}_{t}(D),$ (3.15)
which can be treated as the exact constraint on any solution to QCD, since coming out from the corresponding ST identities, and deriving before going to the formulation of any proper subtraction scheme.This underlines its intrinsic, independent and important status in the theory.
3.1 General solution
The general solution of the exact constraint (3.15) determines the function $\xi=f(q^{2};\xi_{0})$ as follows:
$\xi=f(q^{2};\xi_{0})={\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)}.$ (3.16)
We will call it as the generalized gauge-fixing function approach to QCD as well. It has been exactly (i.e., without making any kind of the simplifications) and uniquely defined in a such new manner for the first time.This expression demonstrates the NL dependence $\xi=f(q^{2};\xi_{0})$ on $\xi_{0}$. The linear relation $\xi\rightarrow\xi_{0}$ will be recovered in the PT $q^{2}\rightarrow\infty$ limit, when the ratio $(\Delta^{2}_{t}(D)/q^{2})$ is to be suppressed in this regime at finite $\xi_{0}$. We have established the general boundary condition in QCD for the behaviour of the gauge-fixing function in the PT limit for the regularized full gluon propagator. Since the general solution (3.16) has been obtained within the ST identities formalism, we will call it as the general solution to QCD. Let us underline that to know it means to know the true gauge structure of the QCD vacuum and its theory at the fundamental quark-gluon level (for more detailed description see section 5 as well).
3.2 Particular solution
The exact constraint (3.15) has also a particular solution, namely
[…]
but these equalities should be put by hand in this case, i.e., they are the prescriptions rather than the exact mathematical results. We will call it as the particular solution to QCD. Let us note in advance that these conditions determine the so-called conventional or, equivalently, conventional PT QCD (though the coupling constant remains always strong apart from the asymptotic free (AF) regime [2, 38, 39]). Precisely, this solution to QCD has been studied the last five decades. In this solution the gauge symmetries of the QCD Lagrangian and its ground state coincide, like in QED (see section 6 as well).
4 Proper subtraction scheme
QCD being much more complicated quantum field gauge theory than QED, has the two satisfied transverse relations (3.14), and so requires much more careful investigation in QCD. They have to be included in a self-consistent way into the proper subtraction scheme. Finally this will lead to the removal of the corresponding QD constants from the theory. How they fixed the gauge-fixing function approach to QCD in the most general and unique ways has been just described here.
The first step in the renormalization program of any gauge theory is the removal of the quadratic UV divergences in order to make the corresponding theory renormalizable in the PT sense. It can be achieved by introducing the proper subtraction scheme in order to separate them from the PT logarithmic divergences. The preliminary step in the regularization program has been already done by introducing the corresponding regulating parameters, mentioned above. Within our approach nothing will depend on how exactly the regulating parameters have been introduced. Even the hard cut-off procedures can be used, since we are mainly interested in the regulating of leading order UV divergences [2]. They have to disappear from the theory after the PT and the non-perturbative (NP) renormalization programs will be performed. Let us now introduce the subtractions for the quark and gluon contributions to the full gluon self-energy as follows:
$\Pi^{s(q)}_{\rho\sigma}(q)=\Pi^{q}_{\rho\sigma}(q)-\Pi^{q}_{\rho\sigma}(0)=\Pi^{q}_{\rho\sigma}(q)-\delta_{\rho\sigma}\Delta^{2}_{q},$ (4.18)
$\Pi^{s(g)}_{\rho\sigma}(q;D)$ $=$ $\Pi^{g}_{\rho\sigma}(q;D)-\Pi^{g}_{\rho\sigma}(0;D)$ (4.19)
$=$ $\Pi^{g}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{g}(D),$
and thus $\Pi^{s(q)}_{\rho\sigma}(0)=\Pi^{s(g)}_{\rho\sigma}(0;D)=0$, by definitions. In these relations $\Delta^{2}_{q}$ is the skeleton loop integral for the quark degrees of freedom at $q=0$, and $\Delta^{2}_{g}(D)$ is the sum of the corresponding skeleton loop integrals at $q=0$ contributing to eq. (2.3), namely
$\Delta^{2}_{g}(D)=\Delta^{2}_{gh}+\Delta^{2}_{1}(D^{2})+\Delta^{2}_{2}(D^{4})+\Delta^{2}_{2^{\prime}}(D^{3}).$ (4.20)
Let us remind that all the QD constants, shown up in eqs. (4.18)-(4.19), as well as the tadpole term $\Delta^{2}_{t}(D)$ itself, are independent from each other and are defined by the skeleton loop integrals, which have been already regularized from above and below. The subtraction at zero is to be understood in a such way that we subtract at the external gluon momentum $q^{2}=-\mu^{2}$ [2] with $\mu^{2}\rightarrow 0$ final limit. However, due to the self-interaction of the massless gluon modes, some of these constants (apart from the tadpole term) may depend on the internal zero gluon momentum (if it is connected to any closed loop, for example such as which can appear in the iteration solution of the gluon SD equation [6]). Then it has to be replaced by the subtraction at $q^{2}_{i}=-M^{2}_{i}$ with $M^{2}_{i}\rightarrow 0$ final limit as well, and subindex $i$ determines the number of such internal gluons. For the any single full gluon propagator the subindex $i$ can be treated as the number of the necessary subtractions made in it in accordance with the rules of the theory of distributions (generalized functions) [26].
Also it is necessary to emphasize that the subtractions (4.17)-(4.18) have not been introduced by hand. They are equivalent to add zero to the corresponding identity. For example $\Pi^{q}_{\rho\sigma}(q)=\Pi^{q}_{\rho\sigma}(q)-\Pi^{q}_{\rho\sigma}(0)+\Pi^{q}_{\rho\sigma}(0)$, so that the initial quark contribution to the gluon self-energy is not changed. It makes it possible to present the initial quark contribution $\Pi^{q}_{\rho\sigma}(q)=\Pi^{s(q)}_{\rho\sigma}(q)+\Pi^{q}_{\rho\sigma}(0)$ as a sum of the two terms, one of which shows up all the corresponding QD but regularized constants. The first term is a regular function of the external gluon momentum $q$, by definition, and it may be only logarithmically divergent at large $q$. In other words, we separate the NP QD constants from the PT logarithmically divergent terms in a gauge-invariant way, that is why our subtraction scheme is called as a proper one. The same is true for the initial gluon contribution as well. Such an exact separation will be very useful for the renormalization programs of any kind to be performed for the single full gluon propagator, and we consider this as an advantage of the formalism developed here. For the interacting full gluon propagators these independent terms will interact with each other, but such a separation will simplify the corresponding renormalization programs, anyway.
The independent tensor decompositions of the quark and gluon degrees of freedom, which appear in the subtraction relations (4.18)-(4.19), are
$\Pi^{s(q)}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(q)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(q)}_{l}(q^{2}),$
$\Pi^{q}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{q}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{q}_{l}(q^{2}),$ (4.21)
and
$\Pi^{s(g)}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(g)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(g)}_{l}(q^{2}),$
$\Pi^{g}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{g}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{g}_{l}(q^{2}),$ (4.22)
respectively. In all the quantities above and below the dependence on $D$ is omitted, for simplicity, and will be restored when necessary. Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary. However, all the invariant functions $\Pi^{s(q)}_{t}(q^{2})$, $\Pi^{s(q)}_{l}(q^{2})$ as well as $\Pi^{s(g)}_{t}(q^{2})$, $\Pi^{s(g)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the initial subtractions (4.18)-(4.19).
Substituting all these decompositions (4.21)-(4.22) into the subtractions (4.18)-(4.19), and doing some tensor algebra derivations, one finally obtains
$\Pi^{s(q)}_{t}(q^{2})$ $=$ $\Pi^{q}_{t}(q^{2})-{\Delta^{2}_{q}\over q^{2}},$
$\Pi^{s(q)}_{l}(q^{2})$ $=$ $\Pi^{q}_{l}(q^{2})+{\Delta^{2}_{q}\over q^{2}},$ (4.23)
and
$\Pi^{s(g)}_{t}(q^{2})$ $=$ $\Pi^{g}_{t}(q^{2})-{\Delta^{2}_{g}(D)\over q^{2}},$
$\Pi^{s(g)}_{l}(q^{2})$ $=$ $\Pi^{g}_{l}(q^{2})+{\Delta^{2}_{g}(D)\over q^{2}}.$ (4.24)
Using these relations it is easy to show that for the regularized (i.e., finite) quark and gluon contributions to the initial gluon self-energy their values at zero always are $\Pi^{q}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{q}$ and $\Pi^{g}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{g}(D)$, indeed, i.e, they are general ones, and have been already used in the subtraction relations (4.18)-(4.19) above, as it has to be. Also these relations will be true if expressed in terms of the above-mentioned subtraction point $q^{2}=-\mu^{2}$.
An interesting observation follows from the relations (4.23) and (4.24), namely $\Pi^{q}_{t}(q^{2})+\Pi^{q}_{l}(q^{2})=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(q)}_{l}(q^{2})$ and $\Pi^{g}_{t}(q^{2})+\Pi^{g}_{l}(q^{2})=\Pi^{s(g)}_{t}(q^{2})+\Pi^{s(g)}_{l}(q^{2})$. So that the corresponding sums do not depend on the corresponding QD but regularized constants, which once more underlines the general character of our subtraction scheme. Also, the sum of the subtracted invariant functions is never zero, though the subtracted contributions to the full gluon self-energy are zero at $q^{2}=0$, by definitions, see relations (4.18)-(4.19).
As we already know, the quark and gluon contributions to the full gluon self-energy are transverse $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$, see the relations (3.14). Then from the second of the relations (4.21) and (4.22) it follows that $\Pi^{q}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})=0$ as well. So that the second relations in (4.23) and (4.24) will be reduced to
$\Pi^{s(q)}_{l}(q^{2})={\Delta^{2}_{q}\over q^{2}},\quad\Pi^{s(g)}_{l}(q^{2};D)={\Delta^{2}_{g}(D)\over q^{2}},$ (4.25)
where we restored the dependence on $D$ in the second relation. However, these relations are impossible, since the corresponding invariant functions cannot have the pole-type singularities, by definitions, as explained above. One has to put these constants to zero on this general mathematical basis, i.e., $\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0$, and thus $\Pi^{s(q)}_{l}(q^{2})=\Pi^{s(g)}_{l}(q^{2};D)=0$ as well. All our results for the QD constants can be now summarized as follows:
$\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0,\quad\Delta^{2}_{t}(D)\neq 0,$ (4.26)
and hence $\Delta^{2}_{gh}=\Delta^{2}_{1}(D^{2})=\Delta^{2}_{2}(D^{4})=\Delta^{2}_{2^{\prime}}(D^{3})=0$, as well because of the relation (4.20). The system of the relations (4.26) is not only a general but it is a unique one as well. They are not prescriptions, since based on a rigorous tensor algebra derivation rules, i.e., they are exact mathematical results. The characteristic feature of this system is that the tadpole term $\Delta_{t}^{2}(D)$ remains intact. The mathematical reason is that for the constant tadpole term such kind of the non-trivial relations (4.18)-(4.19) does not exist (constant minus the same constant is always zero). The essential dynamical source of its survival in the theory is that only the tadpole term can generate a mass squared scale parameter [6]. All the other quark and gluon terms, contributing to the full gluon self-energy, cannot do this, and thus none of their subtracted counterparts with their quark and gluon constants.
All these constants, defined by the corresponding skeleton loop integrals, are QD at the upper limit. They have to be removed/disregarded from the theory on the general mathematical basis, i.e., put zero, as described above, apart from the tadpole term (2.4). In comparison with it, all the other QD quark and gluon constants will be called the tadpole-like/type terms. The general question arises now, namely how to understand these exact equalities to zero in the relations (4.26)? These equalities mean that any tadpole-like term which may appear in the theory by any possible way has to be discarded/disregarded in the theory, i.e, put zero, independently from any other tadpole-like terms. In this way such sums will be always zero, indeed. Such constants may appear even not as the result of the subtractions, but, for example as the result of the NL iteration procedure for the full gluon propagator [6] (see appendix A as well). In this case the tadpole-like terms may be even multiplied by some regularized functions. Whatever their origins would have been in the full gluon propagator, all of them belong to the infinite manifold of the relations (4.26), and thus should be always removed from the theory.
Collecting our results obtained above for the quark and gluon contributions to the gluon self-energy, one obtains
$\Pi^{q}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(q)}_{t}(q^{2}),$
$\Pi^{g}_{\rho\sigma}(q;D)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(g)}_{t}(q^{2};D),$ (4.27)
since they become transverse and coincide with their subtracted counterparts. Reminding also that the invariant functions present in the previous relations are regular functions at small $q^{2}$, i.e., they have no pole-type singularities, and may be only logarithmically divergent at large $q^{2}$. This becomes possible only due to the satisfied transverse relations (3.14) for the quark and gluon degrees of freedom. Just this decreases the quadratic UV divergences of the corresponding skeleton loop integrals to a logarithmic ones, as it has been described in detail in this section.
5 General solution to QCD
Substituting the sum of the terms (4.27) into the eq. (3.12), on account of the relation $\delta_{\rho\sigma}=T_{\rho\sigma}(q)+L_{\rho\sigma}(q)$ and further into the gluon SD eq. (2.5), one finally gets
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{T}_{\mu\nu}(q;D)+D^{L}_{\mu\nu}(q;D),$ (5.28)
where
$D^{T}_{\mu\nu}(q;D)=D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)[q^{2}\Pi(q^{2};D)+\Delta^{2}_{t}(D)]D_{\sigma\nu}(q),$ (5.29)
and
$D^{L}_{\mu\nu}(q;D)=D^{0}_{\mu\rho}(q)iL_{\rho\sigma}(q)\Delta^{2}_{t}(D)D_{\sigma\nu}(q),$ (5.30)
respectively. $\Pi(q^{2};D)=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(g)}_{t}(q^{2};D)$, and this invariant function is regular at zero and may have only the logarithmic divergences in the PT $q^{2}\rightarrow\infty$ limit. The gluon SD eq. (5.28) is now equivalent to the initial gluon SD eq. (2.5), only being re-written in the different form.
Combining the general gluon SD eq. (5.28) with the decompositions (1.2) and (1.3), and on account of the relations (5.29) and (5.30), one obtains
$d(q^{2})={1\over 1+\Pi(q^{2};D)+(\Delta^{2}_{t}(D)/q^{2})}.$ (5.31)
This relation is the NL transcendental equation for the different invariant functions $d(q^{2}),\ \Pi(q^{2};D)$ and the constant $\Delta^{2}_{t}(D)$, i.e., $d=f(D(d))$. Nevertheless, from this expression is clearly seen that in the PT $q^{2}\rightarrow\infty$ regime, the contribution $(\Delta^{2}_{t}(D)/q^{2})$ can be neglected, but the invariant function $\Pi(q^{2};D)$ may still depend on this ratio under the PT logarithms. In the NP region of finite and small gluon momenta this term is dominant, and the dependence of $d(q^{2})$ on $\Delta^{2}_{t}(D)$ may be much more complicated due to the transcendental character of eq (5.31).
Contracting the full gluon SD eq. (5.28), on account of the relation (5.30), with $q_{\mu}$ and $q_{\nu}$, and substituting its result into the general ST identity (3.7), one arrives at
$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi_{0}\left(1-\xi{\Delta^{2}_{t}(D)\over q^{2}}\right)=i\xi,$ (5.32)
which solution is
$\xi\equiv\xi(q^{2};\xi_{0})={\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)},$ (5.33)
i.e., in this case the gauge-fixing parameter becomes the function $\xi\equiv\xi(q^{2};\xi_{0})$ and thus it is not a constant equal to $\xi_{0}$. Let us point out that the expression (5.33) coincides with the general solution (3.16), which is as it should be. So that, its asymptotic properties have been already discussed in the text after eq. (3.16). The generalized gauge-fixing function (5.33) is the known function of its arguments, while the relation (5.31) is the NL transcendental one. Behind the general inequality $\xi\neq\xi_{0}$ is the regularized constant $\Delta^{2}_{t}(D)$ as its dynamical source.
Substituting equations (5.28) and (5.29)-(5.30) into the general decomposition (3.8) for the full gluon propagator, one finally obtains
$D_{\mu\nu}(q)$ $=$ $iT_{\mu\nu}(q){1\over q^{2}+q^{2}\Pi(q^{2};D)+\Delta^{2}_{t}(D)}$ (5.34)
$+$ $iL_{\mu\nu}(q){\lambda^{-1}\over q^{2}+\lambda^{-1}\Delta^{2}_{t}(D)},$
where we have introduced the useful notation $\xi_{0}=\lambda^{-1}$ [4]. The corresponding ST identity now becomes
$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi(q^{2};\lambda^{-1})=i{\lambda^{-1}q^{2}\over q^{2}+\lambda^{-1}\Delta^{2}_{t}(D)}.$ (5.35)
This is the generalized gauge-fixing function since it depends on the constant $\Delta^{2}_{t}(D)$, and when it is zero, one recovers the gauge-fixing parameter for the free gluon propagator. In this ST identity the gauge-fixing parameter $\lambda^{-1}$ is convenient to vary continuously from zero to infinity. The functional dependence of the generalized gauge-fixing function $\xi$ (5.33) is fixed up to an arbitrary gauge-fixing parameter $\xi_{0}=\lambda^{-1}$. Unless we fix it, and thus $\xi$ itself, we will call such situation as the general gauge dependence (GGD), see equations (3.7)-(3.8), (5.32)-(5.34). Choosing $\xi_{0}=\lambda^{-1}$ explicitly, we will call such situation as the explicit gauge dependence (EGD). For example, $\xi_{0}=\lambda^{-1}=0$ is called the unitary (Landau) gauge, $\xi_{0}=\lambda^{-1}=1$ is called the t’ Hooft–Feynman gauge, etc. [30, 31, 32, 33]. The formal $\xi_{0}=\lambda^{-1}=\infty$ limit is called as the canonical gauge in [32]. This distinction seems a mere convention, but, nevertheless, it is useful one in QCD because of the presence of a mass scale parameter in its ground state. The generalized gauge requires that there is no other functional expression for $\xi$, apart from given by the relation (5.32) at finite $\xi_{0}=\lambda^{-1}$, in the full gluon propagator (5.34) and the ST identity (5.35) for the regularized gluon fields.
The system of the regularized equations (5.34)-(5.35) constitutes that the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is not a symmetry of its ground state, and it has been exactly and uniquely defined. They are present in the form suitable for the NP renormalization program to be performed. The tadpole term enters the full gluon self-energy linearly, see the expression (3.12). However, in the full gluon propagator (5.34) it appears in the NL way, because its contribution has been summed up with the help of the gluon SD eq. (5.28). The term $(\Delta^{2}_{t}(D)/q^{2})$ in the expressions (5.34)-(5.34) is to be suppressed in the PT $q^{2}\rightarrow\infty$ at the finite that $\xi_{0}=\lambda^{-1}$ (the above-mentioned canonical gauge will be investigated in detail in the forthcoming paper). The full gluon propagator will behave like the free gluon propagator $\sim 1/q^{2}$. This is the one of the necessary constraints that a theory is perturbative renormalizable. Other ones such as the corresponding behaviour of the spinor Green’s function, a unitary of $S$-matrix and analyticity (causality) [2], evidently are beyond the scope of the present work. Here we have proven that the true dynamical and gauge structures of the QCD/YM ground state are much more complicated than it is required by its Lagrangian’s gauge symmetry. But despite the gauge symmetry is broken in the QCD ground state, its PT renormalizability is not affected within the QCD full gluon propagator in the generalized gauge (5.33), i.e., it has the PT renormalizable behaviour at large $q^{2}$.
Concluding, it is worth noting that in terms of the transverse relations and the corresponding QD regularized constants as well as the corresponding relations between $\xi$ and $\xi_{0}$, the general solution can be shown explicitly as the following system of the relations, namely
$q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)$ $=$ $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0,$
$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)$ $=$ $q^{2}\Delta^{2}_{t}(D),$ (5.36)
and
$\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0,\ \Delta^{2}_{t}(D)\neq 0,\ \xi\neq\xi_{0}=\lambda^{-1}.$ (5.37)
It is necessary to underline that none of the prescriptions (i.e., putting some of them zero by hand) have been introduced and none of the truncations, approximations and assumptions have been done as well in the obtaining of these relations. In other words, all of them are exact mathematical results in the general solution to QCD.
6 Particular solution to QCD
The widely used up to present days the linear relation $\xi=\xi_{0}$ is only a particular solution of the derived here the exact constraint (3.15). As mentioned above in section 3, it is defined when the tadpole term has to be put zero by hand, i.e., $\Delta^{2}_{t}(D)=0$, then $\xi=\xi_{0}=\lambda^{-1}$ as well and vice versa. The corresponding expressions for the full gluon propagator and the ST identity in the particular solution to QCD can be independently derived absolutely in the same way as has been derived the corresponding expressions for the general solution to QCD in the previous section. However, the simplest way to show the particular solution is to put formally $\Delta^{2}_{t}(D)=0$ in the system of eqs. (5.34)-(5.35), then it looks like
$D^{PT}_{\mu\nu}(q)={iT_{\mu\nu}(q)\over q^{2}[1+\Pi(q^{2};D^{PT})]}+i\lambda^{-1}L_{\mu\nu}(q){1\over q^{2}},$ (6.38)
so that the gluon invariant function is $d_{PT}(q^{2})=[1+\Pi(q^{2};D^{PT})]^{-1}$ and formally neglecting the invariant function $\Pi(q^{2};D^{PT})$, i.e., putting it zero by hand $\Pi(q^{2};D^{PT})=0$, one obtains the free gluon propagator (3.9). The ST identities in this case are
$q_{\mu}q_{\nu}D^{PT}_{\mu\nu}(q)=q_{\mu}q_{\nu}D^{0}_{\mu\nu}(q)=i\lambda^{-1}.$ (6.39)
The corresponding gluon SD eq. (5.28) now becomes
$D^{PT}_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{T}_{\mu\rho}(q),$ (6.40)
where
$D^{T}_{\mu\nu}(q)=D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)q^{2}\Pi(q^{2};D^{PT})D^{PT}_{\sigma\nu}(q).$ (6.41)
From above and now on in this section we denote the corresponding full gluon propagator $D$ by $D^{PT}$ (for the explanation see the text below). Obviously, the expression for the PT full gluon propagator (6.38) describes the propagation of the PT massless gluons, since it has the PT singularity on the mass-shell $q^{2}=0$, i.e., the singularity of the free gluon propagator $\sim 1/q^{2}$. Let us note in advance that the equality $\xi=\xi_{0}=\lambda^{-1}$ takes place only for the regularized massless gluon fields. For their renormalized counterparts these gauge-fixing parameters differ by the corresponding renormalization constant.
The system of the regularized eqs.(6.38)-(6.41), which is uniquely defined, is free of all the types of the scale parameters having the dimensions of a mass squared, forbidden by the exact gauge symmetry of the QCD Lagrangian. Therefore, it constitutes that the gauge symmetry of the QCD ground state, reflected by this system, coincides with the symmetry of its Lagrangian. Such coincidence is analogous to QED where the abelian $U(1)$ gauge symmetry of the Lagrangian is the same as of its ground state (omitting formally the YM degrees of freedom here, this coincidence has been proven just in this way in [10]). However, it is necessary to remind that the coupling constant remains strong, so to call it as $D^{PT}$ is only a mere convention, that is why the particular solution to QCD can be called as conventional QCD as well. Just this conventional solution to QCD has been investigated in full details the last five decades since its formulation in [1].
The essential feature of the above-mentioned coincidence is the equality $\xi=\xi_{0}$. In this connection, let us note that one can start from the expression (3.8). It is the general decomposition of the full gluon propagator, depending on the one variable, into the independent tensor structures. Then it will automatically satisfy to the relation (3.7), called the ST identity, determining thus the gauge-fixing parameter. This has nothing to do with the statement that the ST identity is a consequence of the exact gauge symmetry. It becomes its consequence, indeed, when $\xi$ is fixed by the relation $\xi=\xi_{0}=\lambda^{-1}$. The well-known PT linear relation $\xi=\xi_{0}$ is the consequence of the exact gauge invariance both for the QCD Lagrangian and its ground state. In principle, the PT QCD or, equivalently, the conventional QCD is a well-elaborated quantum gauge field theory [2, 3, 4, 5, 6, 7, 8, 9], and for its framework of the corresponding dynamical equations of motion see recent review [34] (and references therein). However, this theory is not our concern (but for some clarifying discussion see appendix B). We present it here within our formalism for the readers convenience to directly compare it with the general solution to QCD.
For this aim it is useful to reproduce the system of the relations (5.36)-(5.37) for the PT QCD. It will look like
$q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)$ $=$ $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D^{PT})=0,$
$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D^{PT})$ $=$ $0,$ (6.42)
and
$\Delta^{2}_{q}=\Delta^{2}_{g}(D^{PT})=0,\ \Delta^{2}_{t}(D^{PT})=0,\ \xi=\xi_{0}=\lambda^{-1}.$ (6.43)
The first equalities in these relations coincide with the corresponding equalities in the relations (5.36)-(5.37), i.e., in both theories the general and the PT QCD or, equivalently, the particular QCD, they are exact mathematical results. At the same time, the second equalities in these relations remain prescriptions, since they have to be put zero by hand, as emphasized above.
The dimensional regularization method (DRM) [4, 6, 27, 28, 29], formulated within the framework of the conventional QCD, provides a gauge-invariant scheme to correctly calculate the finite parts of the generally QD loop integrals, while omitting their QD constants, i.e., simply to ignore them and to deal further only with the logarithmic divergences of the PT. This was a prescription rather than an exact result, as pointed out in [4]. However, from now on this prescription has been put on a firm mathematical ground. How to understand such kind of equalities has been explained in section 5.
7 Discussion
Let us now discuss some interesting physical consequences of the general solution to QCD. It requires the explicit presence of the tadpole term in the QCD ground state at the fundamental quark-gluon level, i.e., it cannot be discarded in the QCD/YM theory by any means. This was exactly proven in section 4 and used in section 5. Then it will drastically change the structure of the full gluon propagator in the deep IR region $q^{2}\rightarrow 0$. For the sake of transparency and clarity let us use the free gluon propagator in the t’ Hooft-Feynman gauge $\lambda^{-1}=1$, i.e., $D^{0}_{\mu\nu}(q)=i{\delta_{\mu\nu}/q^{2}}$. Then the full gluon propagator (5.28), on account of the definitions (5.29) and (5.30), up to the first iteration term looks like
$D_{\mu\nu}(q)\sim i{\delta_{\mu\nu}\over q^{2}}-i{T_{\mu\nu}(q)\over q^{2}}\Pi(q^{2};D)-iT_{\mu\nu}(q){\Delta^{2}_{t}(D)\over(q^{2})^{2}}+...\ ,$ (7.44)
where the invariant function $\Pi(q^{2};D)$ is a regular one of its argument, while all the tadpole-like terms up to one loop approximation have been omitted due to the relations (4.26). See appendix A as well, where this iteration is discussed in more detail. Also here and below not shown the terms proportional to the longitudinal component $L_{\mu\nu}(q)$ since finally they will be summed up to the known function, see expression (5.33). The higher-order iteration terms will be much more singular with respect to the general ratio $(\Delta^{2}_{t}(D)/q^{2})$, since each iteration invokes additional powers of this ratio. Because of the NL character of the iteration procedure any next iteration will contribute into the previous iteration terms, and thus they become the sums of the infinite number of terms [35], namely
$D_{\mu\nu}(q)\sim iT_{\mu\nu}{\Delta^{2}_{t}(D)\over(q^{2})^{2}}\sum_{k=0}^{\infty}\Bigl{(}{\Delta^{2}_{t}(D)\over q^{2}}\Bigr{)}^{k}\Phi_{k}(g^{2},...)+...,$ (7.45)
where the terms of order $O_{\mu\nu}(1/q^{2})$ are not shown, for simplicity. In this iteration expansion is explicitly shown only singular part of the corresponding Laurent series, and thus it represents the summation of all the severe IR singularities (more singular than the PT IR singularity $\sim 1/q^{2}$), which can be accumulated by the full gluon propagator. Evidently, such a summation is only possible due to the existence of the tadpole term. The dependence of the coefficients $\Phi_{k}(g^{2},...)$ on the coupling constant is not known, i.e., such kind of series are the cluster expansions, mentioned above [23]. These coefficients, being the sum of the infinite number of terms by themselves, are resulted from the skeleton loop integrals, contributing into full gluon self-energy through the invariant function $\Pi(q^{2};D)$ because of the NL character of the iteration procedure. Evidently, the iteration ab  infinity leads to the essential singularity at $q^{2}\rightarrow 0$! in the full gluon propagator (7.44), which will remain massless, but depending on the tadpole term. Let us note that the same series can be obtained by expanding the initial NL transcendental relation (5.30) in powers of the ratio $(\Delta^{2}_{t}(D)/q^{2})$.
From the NL iteration expressions (7.44)-(7.45) it follows that the full gluon propagator in the IR region ($q^{2}\rightarrow 0$) is dominated by the tadpole contribution, while in the UV region ($q^{2}\rightarrow\infty$) it will be dominated by the free gluon one. This is completely different to QED, where the dressed photon propagator behaves at small photon momentum as its free counterpart ($\sim 1/q^{2}$). This is also true for the full gluon propagator in the PT QCD with unbroken gauge symmetry, investigated in section 6. The principal difference between these theories is due to the existence of the tadpole term in the QCD ground state, which precisely controls the structure of a single full gluon propagator at small $q^{2}$. That is why the dressed PT gluon (6.38) and free gluon states can not appear in the physical spectrum at large distances ($q^{2}\rightarrow 0$), i.e., they will be suppressed in this region by the tadpole term. If this statement will survive the corresponding renormalization program beyond the PT then confinement of the PT and free gluon states will be explained. This is important since there is no any other dynamical mechanism to confine these gluon states to the vacuum of QCD or to keep them inside hadrons. To prove the confinement of the singular full gluon propagator itself due to the above-mentioned essential singularity, though possible, but requires completely different investigation within the theory of distributions [26], and thus is the subject of the future work. Let us note that by the NP program we mean that the full gluon propagator will become free from all the non-physical parameters of the theory in any of its possible solutions (not only in the iteration one).
Eq. (5.34) shows that gluon may acquire mass dynamically if the denominator of this equation becomes zero at some point $q^{2}=M^{2}$. The expression for the free massive vector particle propagator follows in a rather simple way, indeed. For this goal it is necessary to omit the corresponding invariant function and identify the tadpole term with the gluon mass $M^{2}$. In Minkowski metric (by replacing $q^{2}\rightarrow-q^{2}$ and $\delta_{\mu\nu}\rightarrow g_{\mu\nu}$), one gets
$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2})}\left[g_{\mu\nu}-{(1-\lambda^{-1})q_{\mu}q_{\nu}\over(q^{2}-\lambda^{-1}M^{2})}\right],$ (7.46)
i.e., it will be expressed in the Stueckelberg gauge [4]. Contracting this expression with respect to $q_{\mu}$ and $q_{\nu}$, one obtains the ST identity (5.35) by replacing there $q^{2}\rightarrow-q^{2}$. In other words, the Stueckelberg gauge (7.46) is a particular case of the generalized gauge-fixing function (5.33). The PT QCD full gluon propagator (6.38) cannot generate the gluon mass at some point, since it remains always massless. The full gluon propagator (5.34) provides a rather convenient framework in order to formulate and further develop the NP renormalization program for the massive gluon fields. This program will suppose to show that the full gluon propagator (5.34), and hence its free massive counterpart (7.46), can not be put on the mass-shell $q^{2}=M^{2}_{g}$, where $M^{2}_{g}$ is the finite gluon pole mass. Acquiring mass dynamically, the gluon, nevertheless, will remain the off-mass-shell object. So that, it will not appear in the physical spectrum. Some aspects of the massive gluon field’s renormalization are discussed in appendix C.
In the presence of the mass squared scale parameter - the tadpole term - the role of the QCD coupling constant $g^{2}$ becomes unimportant. This is also evidence of the ’dimensional transmutation’, $g^{2}\rightarrow\Delta^{2}_{t}(D)$ [2, 36, 37], which occurs whenever a massless theory acquires mass dynamically. It is a general feature of spontaneously symmetry breaking in field theories. We distinguish between the PT/particular/conventional and general solutions to QCD by the explicit presence of the tadpole term in the latter one, and not by the magnitude of the coupling constant even at the regularized gluon fields level yet. In the both cases the gluon fields remain strongly interacted, apart from the AF regime [38, 39]. So that, the general solution to QCD, investigated in section 5 is possible to call as the NP QCD as well.
Any deviation of the full gluon propagator from the free one in QCD requires the presence of a mass squared scale parameter. Even in the AF regime $q^{2}\rightarrow\infty$ there is a scale breaking,
$D_{\mu\nu}(q)\sim\delta_{\mu\nu}\Bigl{[}{ig^{2}\over 1+g^{2}b_{0}\ln(q^{2}/\Lambda^{2}_{QCD})}\Bigr{]}(1/q^{2}),$ (7.47)
which has been discovered many years ago [37, 38]. When $q^{2}\rightarrow\infty$ the dependence on the coupling constant $g^{2}$ vanishes, indeed. The well-known expression (7.47) is written down in the t’ Hooft-Feynman gauge and in the approximation of the main PT logarithms with the colour group factor $b_{0}>0$. $\Lambda^{2}_{QCD}$ determines the scale of the non-trivial PT dynamics in the QCD. However, in section 6, we have already shown that the PT cannot generate the scale parameter, having the dimension of a mass squared. From the renormalization group equations analysis, it follows that any mass to which can be assigned any physical meaning is exponentially suppressed in the weak coupling limit to every order of the PT [2, 36], namely
$M\sim\mu\exp(-1/b_{0}g^{2}),\quad\quad g^{2}\rightarrow 0,$ (7.48)
where $\mu$ is the arbitrary renormalization point. None a finite mass can survive in the PT weak coupling limit or, equivalently, in the PT $q^{2}\rightarrow\infty$ regime, indeed. So the question where the finite mass comes from? cannot be answered by the PT! It is entirely of the NP origin! It has to come from the IR region which is controlled by the tadpole term, as discussed just above. Its finite version should be somehow related to $\Lambda^{2}_{QCD}$ (reminding that the ratio $(\Delta^{2}_{t}(D)/q^{2})$ may appear under the PT logarithms). The explanation of the AF phenomenon (i.e., a scale breaking in the PT regime) will be given when the NP renormalization program of the singular gluon field configurations will be performed.
8 Conclusions
We have derived the novel exact constraint (3.15) on any solution to QCD. For this aim the ST identities for the gauge particles Green’s functions (propagators), complemented by the satisfied transverse relations for the quark and gluon contributions into the full gluon self-energy, have been used. The first of its solutions is called the general solution to QCD and given by the expression (3.16), which formulates the generalized gauge-fixing function approach to QCD. The existence of a new general solution to QCD makes it possible to prove in the most general and unique mathematical ways that the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is broken in its ground state. However, the PT renormalizability will not be affected within the framework of the full gluon propagator. In the PT $q^{2}\rightarrow\infty$ regime it will have the PT renormalizable behaviour $\sim 1/q^{2}$.
The dynamical source of the Lagrangian’s gauge symmetry breakdown in the ground state has to be identified with the constant tadpole term, having the dimension of a mass squared. It is dynamically generated by the self-interaction of the multiple massless gluon modes involving only the point-like 4-gluon vertex, see eq. (2.4) and Fig. 1. The full 4-gluon vertex survives when all the gluon momenta involved go to zero, becoming thus its point-like counterpart, while the full 3-gluon vertex does not survive in this limit. Just this shows the importance of the point-like 4-gluon vertex in the deep IR structure [10, 23, 40] of the theory. By the design the tadpole term is present in the QCD ground state from the very beginning, i.e., not introducing by hand. It is only one which explicitly contributes into the skeleton loops decomposition of the full gluon propagator (shown in Fig. 1), while all the other tadpole-like terms don’t. Unlike them, being the QD but already regularized constant, the tadpole term is not connected to any external gluon momentum, and thus may generate a mass squared scale parameter [6], indeed. Thus, there is no need to extract such a scale parameter in the QCD vacuum in a some other way. In [10] the importance of the tadpole term has been recognized and exploited on an intuitive basis rather than a logical one. However, from now on its existence cannot be undermined. Besides its combinatorial (mathematical) meaning it has a deep dynamical one as a source of the gauge symmetry violation in the QCD vacuum. Its renormalized (i.e., finite) version will determine a scale of the NP dynamics in QCD and in its true ground state.
Due to these impacts our advance findings are:
a. The novel general constraint (3.15) on any solution to QCD has been exactly derived before any proper subtraction scheme has been introduced. The importance of the ST identities and the corresponding satisfied transverse relations has been explicitly shown.
b. Its exact solution (3.16) determines the generalized gauge-fixing function approach to QCD.
c. The proper subtraction scheme has been formulated to remove from the theory the QD constants, apart from the tadpole term, on the general mathematical basis.
d. The system of the relations (5.36) are exact mathematical results, i.e., they are not prescriptions.
e. The new NP approach to the YM theory and its ground state has been formulated, based on the regularized system of eqs.  (5.34)-(5.35). It is defined up to the invariant function with the known asymptotic properties.
f. This system has a correct the PT limit, and is suitable for the NP renormalization program to be performed. It is called as the general solution to QCD, and just it resolves the dilemma formulated at the end of section 2.
In obtaining all these results no any truncations, approximations and assumptions have been made in dealing with the regularized skeleton loop integrals, contributing to the gluon self-energy. The exact analytic derivations have been done, based on a tensor algebra rigorous rules.
The second of the above-mentioned solutions is called the particular solution to QCD, and determined by the relations $\Delta^{2}_{t}(D)=0$ and $\xi=\xi_{0}$ in the exact constraint (3.15). This solution is nothing else but the well-known PT or, equivalently, conventional QCD, which has been studied last five decades. We confirm that the second solution preserves the gauge symmetry of the QCD Lagrangian in its ground state. Both equalities, shown above, are simply prescriptions, i.e., they have to be put valid by hand, as it has been discussed in section 6. In this connection, let us remind that for the tadpole term does not exist the independent mathematical relation according to which it has to be omitted from the theory, while such kind of the relations exist for the the quark and gluon QD regularized constants (see section 4). That is why the tadpole term has to be put zero by hand for the particular solution, which comes from the solution of the exact constraint (3.15), and see also discussion in appendix B.
We can registered the existence of the two different types of the gluon field configurations in the QCD vacuum: the general one and its particular counterpart with the gauge symmetry broken and unbroken, respectively.
The just derived general solution to QCD with a gauge symmetry breaking, but maintaining its PT renormalizability, will apparently satisfy the boundary conditions in QCD at the fundamental quark-gluon level: the confinement at large distances ($q^{2}\rightarrow 0$) and the AF at short distances ($q^{2}\rightarrow\infty$), as discussed in section 7. The particular or, equivalently, the PT/conventional QCD solution with a gauge symmetry preservation cannot satisfy the boundary condition at large distances. It shows the agreement with the AF regime at short distances, but cannot explain it, see discussion in section 7 as well.
The perspectives for further work are as follows: to formulate the NP multiplicative (MP) renormalization programs for the regularized full gluon propagator (5.34). This is not a simple task, since such kind of a programs have to be different for the massive and severely singular gluon field configurations, briefly discussed in section 7. For this purpose it is necessary to take into account the transcendental and NL structure of eq. (5.31). They are also needed to be formulated in the most general and unique ways. To ensure a unitary of $S$-matrix elements by these NP MP renormalization programs should be requested as well. In general, our results open a new wide research areas in modern physics, leading to the multiple theoretical and phenomenological applications and comparing them with the results of the conventional QCD.
To understand a mass dynamical generation at the quark-gluon level is the first necessary step to understand the existence of the physical mass spectrum at the hadronic level in QCD as a theory of strong interactions. The confinement first and only after the PCAC phase transitions in QCD have to be understood [2, 23]. QCD is a self-consistent quantum gauge field theory. It does not need any extra degrees of freedom (such as Higgs fields [41]) to generate a mass at the fundamental quark-gluon level.
Concluding, a few remarks are in order. The Elitzur’s theorem [42] has been proven in lattice field theory for the abelian system like QED (see appendix D as well). To the non-abelian systems like YM fields, it also can be applied, but only within its particular solution, as discussed above. At the same time, this theorem cannot be applied to the YM fields in the frame work of its exact solution for the generalized gauge-fixing function in the full gluon propagator. In this paper we have just shown that the gauge symmetry of the continuos YM theory can be broken in its ground state. Behind this effect is the mass squared parameter, which is nothing else but the renormalized tadpole term. It survives the NP renormalization but does not affect the PT renormalizability, as underlined just above. The relevant existing knowledge on a mass dynamical generation and spontaneously symmetry breaking subjects has been taken into account in the list of publications, presented in our paper (and see references in each of the cited items there). Also, let us note that contrary to the ghost term, the QCD Lagrangian’s gauge-fixing term is not explicitly present in the full gluon self-energy, that is why it is left out of our attention. It would be interesting to establish the character of its possible connection (if any) to the NL generalized gauge-fixing function approach. Apparently, this might be done only after the corresponding NP MP renormalization programs will be performed, not before, in our opinion.
9 Summary
Briefly, we can summarize our results as follows:
The exact derivation of the general constraint (3.15) on any solution to QCD is present. This has been done in the mathematically understandable way for a broad readership of researches, since exploiting only the tensor algebra derivations rigorous rules, widely used in the theoretical physics, as pointed out above. It has the two possible exact solutions, in principle, independent from each other.
The general solution to QCD is the solution with the gauge symmetry broken in its ground state. The dynamical source of this groundbreaking effect is the tadpole term. The general QCD remains the PT renormalizable, nevertheless. It will satisfy the boundary conditions at the fundamental quark-gluon level, as discussed in section 7.
The particular or, equivalently, the PT/conventional solution to QCD is the solution with unbroken gauge symmetry in its ground state. It is the PT renormalizable as well, but not satisfying the boundary condition at large distances, as discussed in section 7. Its gauge symmetry hides the true role of the tadpole term in the QCD vacuum.
Acknowledgments
The authors are grateful to P. Levai, P. Forgács, J. Nyiri, T.S. Biró, Gy. Kluge, M. Vasúth, P. Ván, G. Somogyi, S. Van Thurenhout for useful suggestions, remarks, discussions and help. The work was supported by the Hungarian National Research, Development and Innovation Office (NKFIH) under the contract numbers OTKA K135515 and NKFIH 2019-2.1.11-TÉT-2019-00078, 2019-2.1.11-TÉT-2019-00050, the Wigner Scientific Computing Laboratory.
Appendix A The full gluon propagator up to one loop
Figure 2: The SD equation for the full gluon propagator up to one loop as present in [6]. It is useful to compare it with Fig. 1 and its description in the text after it.
In order to explicitly compare the tadpole term (2.4) with the other tadpole-type terms let us consider the iteration of the gluon SD eq. (2.1), shown in Fig. 1, up to one loop, shown in Fig. A.2. The last two terms in this figure appear because of the NL nature of the iteration series, when the next iterations contribute into the previous ones. That is the reason why such multiplied tadpole-type terms are not present in the skeleton loop decomposition for the full gluon propagator in Fig. 1, but have to be shown up in the iteration terms, seen in Fig. A.2. This iteration has been exactly calculated in [6] by using the DRM, while rightfully now ignoring the tadpole term itself and all the other tadpole-likes ones, in accordance with the relations (6.43). It is worth emphasizing that in [6] this iteration has not been called the PT expansion up to $O(g^{2})$, though looks like it, but expansion up to $O(\hbar)$. It is well-known that in QCD the expansion in powers of the coupling constant does not make any sense because it is strong. The PT expansion in QCD makes sense only in the AF [37, 38] regime when it becomes weak. In order to perform analytical derivations below, we will take the interaction vertices as the point-like ones. The external gluon propagators will be taken as the free ones. All the skeleton loop propagators will remain as the full ones. Then the corresponding iteration will look like Fig. A.2, and hence can be called as quasi-one loop expansion, discussed in some details in section 7, when it was called as the cluster one. Note that below we do not show explicitly the dependence on the coupling constant, since our skeleton loop quark, gluon and ghost propagators are the full ones, so that the exact dependence on the coupling constant of the expressions below is not known. Also we will omit below all the colour group factors and other overall numerical numbers, since all of them are unimportant here.
However, it is instructive first to show explicitly how works our subtraction scheme, using only the third diagram in the second line of Fig.A.2, for simplicity. From [6] by replacement $\mu\rightarrow\rho$ and $\nu\rightarrow\sigma$ and fixing $\omega=2$ and $\delta_{\rho\rho}=4$ in eq. (6.22) there, one finally obtains an analytical expression for this term as follows:
$\Pi^{(3)}_{\rho\sigma}(q)\sim\int d^{4}l\ {N_{\rho\sigma}(l,q)\over l^{2}(l+q)^{2}},$ (1.49)
where, apart from below, we put the gluon loop propagators as free ones in the t’ Hooft-Feynman gauge as well, for the sake of transparency of our derivations for this diagram. The tensor function $N_{\rho\sigma}(l,q)$ is
$N_{\rho\sigma}(l,q)$ $=$ $10l_{\rho}l_{\sigma}+5(l_{\rho}q_{\sigma}+q_{\rho}l_{\sigma})-2q_{\rho}q_{\sigma}$ (1.50)
$+$ $(5q^{2}+2(lq)+2l^{2})\delta_{\rho\sigma}.$
Putting in eqs. (A.1) and (A.2) $q=0$, one gets
$\Pi^{(3)}_{\rho\sigma}(0)\sim\int d^{4}l\ {N_{\rho\sigma}(l,0)\over(l^{2})^{2}}=\int d^{4}l\ {(10l_{\rho}l_{\sigma}+2l^{2}\delta_{\rho\sigma})\over(l^{2})^{2}},$ (1.51)
and, obviously, this integral is quadratically divergent at the upper limit, but is assumed to be regularized from above and below from the very beginning, as emphasized throughout this work.
Let us now compose the subtracted counterpart of the initial integral (A.1), in accordance with the subtraction (4.18), as follows:
$\Pi^{(3s)}_{\rho\sigma}(q)=\Pi^{(3)}_{\rho\sigma}(q)-\Pi^{(3)}_{\rho\sigma}(0)\sim\int d^{4}l\ {N^{s}_{\rho\sigma}(l,q)\over(l^{2})^{2}(l+q)^{2}},$ (1.52)
where
$N^{s}_{\rho\sigma}(l,q)$ $=$ $5l^{2}(l_{\rho}q_{\sigma}+q_{\rho}l_{\sigma})-20(ql)l_{\rho}l_{\sigma}-2l^{2}q_{\rho}q_{\sigma}$ (1.53)
$+$ $l^{2}(q^{2}-2(lq))\delta_{\rho\sigma}-10q^{2}l_{\rho}l_{\sigma}.$
[…]
The tadpole term up to one skeleton loop, shown first in the third line of Fig. A.2, analytically can be written down as follows:
$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lT^{0}_{\rho\sigma\alpha\beta}D_{\alpha\beta}(l)=\delta_{\rho\sigma}\Delta^{2}_{t}(D),$ (1.54)
see eq. (2.4) as well. Being the QD but already regularized constant, it is not connected to any external gluon momentum. In this respect it is different from all other tadpole-like terms which appear as a result of the subtraction scheme or in the NL iteration procedure. Also it directly generate the mass squared parameter, and thus explicitly violets the gauge symmetry of the QCD Lagrangian in its ground state [6]. We have shown that in the PT QCD or,equivalently, particular QCD (described in section 6) it was a prescription to neglect it, but in the full QCD or, equivalently, NP QCD (investigated in sections 4 and 5) it should remain intact. At the time, all other tadpole-type terms have to be disregarded on the exact mathematical basis in both theories.
Let us now investigate the tadpole-like term with the gluon skeleton loop, shown second in the third line of Fig. A.2. Its analytical expression is
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l\ T^{0}_{\nu\zeta\mu}(l,p)D_{\nu\zeta}(l),$ (1.55)
where the gluon momentum $p$ is, in fact, zero, i.e., $p=0$ but it is convenient to go to this limit at the final step only. Both gluon propagators in the t’ Hooft-Feynman gauge are $D^{0}_{\mu\mu^{\prime}}(p)=\delta_{\mu\mu^{\prime}}/p^{2}$ and $D_{\nu\zeta}(l)=\delta_{\nu\zeta}d(l^{2})/l^{2}$, where $d(l^{2})$ is the corresponding invariant function of the full gluon propagator. Let us remind that such kind of the skeleton loop integrals are assumed to be regularized from above and below. Using now the Euclidean space Feynman rules for the corresponding vertices present in [6], one arrives at
$T^{0}_{\rho\sigma\mu^{\prime}}(q,p)=-(2q+p)_{\mu^{\prime}}\delta_{\rho\sigma}+(q-p)_{\sigma}\delta_{\rho\mu^{\prime}}+(2p+q)_{\rho}\delta_{\sigma\mu^{\prime}}$ (1.56)
and
$T^{0}_{\nu\zeta\mu}(l,p)=2(l+p)_{\mu}\delta_{\nu\zeta}-l_{\zeta}\delta_{\nu\mu}-(2p+l)_{\nu}\delta_{\zeta\mu}.$ (1.57)
Substituting these expressions into the eq. (A.7), one finally obtains
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p){2\over p^{2}}\int d^{4}l{d(l^{2})\over l^{2}}[3l_{\mu}+p_{\mu}].$ (1.58)
By the symmetry integration ($d^{4}l=l^{3}dl=(1/2)l^{2}dl^{2}$ and all the overall numerical numbers due to the integration over angular variables in four-dimensional Euclidean space [6] omitted) the first loop integral is zero, while the second one leads to
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p)p_{\mu}{2\over p^{2}}\times\tilde{\Delta}^{2}_{g}(d),$ (1.59)
where $\tilde{\Delta}^{2}_{g}(d)=\int d^{4}l(d(l^{2})/l^{2})$ is the QD but regularized skeleton loop integral, having the dimension of a mass squared. It is one of the constants present in the relations (4.26). It should be discarded on the general mathematical basis within our approach, i.e., put zero, so that $\tilde{\Delta}^{2}_{g}(d)=0$, even before the final $p^{2}=-M^{2}\rightarrow 0$ limit. Let us note that the cancellation of the pole $1/p^{2}=-1/M^{2}$ in eq. (A.11) can be always achieved by going to the dimensionless variable $x=l/M$ in the skeleton loop integral (A.10), i.e., this pole is not a problem here. Thus, $\Pi^{(4)}_{\rho\sigma}(q,p)=0$.
Let us finally investigate the tadpole-like term with the ghost skeleton loop, shown third in the third line of Fig. A.2. In analogy with eq. (A.7) its analytical expression is
$\Pi^{(5)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l{l_{\mu}\over l^{2}}d_{gh}(l^{2}),$ (1.60)
where $d_{gh}(l^{2})$ is the invariant function for the full ghost propagator $G(l)$. The vertex $T^{0}_{\rho\sigma\mu^{\prime}}(q,p)$ is explicitly shown in the relations (A.8), and the remaining free gluon propagator is also the same. The altered rule [6] for the gluon-ghost vertex gives $(1/2)[(l_{\mu}+l{\mu})=l_{\mu}$ and the gluon momentum $p$ is zero again at the final stage. Since the nominator of this skeleton loop integral contains the loop variable linearly, then by the symmetry integration it is simply zero from the very beginning, i.e., $\Pi^{(5)}_{\rho\sigma}(q,p)=0$.
Concluding, neither the tadpole-like term with the gluon skeleton loop nor its ghost counterpart contribute into the full gluon propagator. In these cases the subtractions with respect to the external gluon momentum $q$ (though possible), but were not even necessary to perform, since the final results would have been the same zeros. The primary goal of this work is to prove that the tadpole term cannot be disregarded by any means. All the other tadpole-like terms are to be always discarded on the general basis, i.e., put zeros, as explicitly demonstrated here as well. This was not a prescription, but being an exact mathematical result within our formalism.
Appendix B The origin of the conventional QCD
In order to find the general solution to QCD, we present the full gluon self-energy (2.2) as a sum of the three independent terms, i.e., we distinguish between the gluon part and the tadpole term, though their sum forms the YM part of QCD, since the both are determined by the different self-interaction of multiple massless gluon modes. However, not doing so, one needs to present the full gluon self-energy in terms of the quark and YM compositions, i.e., one has to operate with
$\Pi_{\rho\sigma}(q;D)=\Pi^{q}_{\rho\sigma}(q)+\Pi^{YM}_{\rho\sigma}(q;D),$ (2.61)
where
$\Pi^{YM}_{\rho\sigma}(q;D)=\Pi^{g}_{\rho\sigma}(q;D)+\delta_{\rho\sigma}\Delta^{2}_{t}(D).$ (2.62)
The relation (3.13) in this case becomes
$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{YM}_{\rho\sigma}(q;D)]={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (2.63)
As we already know, the quark contribution is always satisfied, i.e., $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=0$, which leads to $\Delta^{2}_{q}=0$. Thus, the relation (B.3) is reduced to
$q_{\rho}q_{\sigma}\Pi^{YM}_{\rho\sigma}(q;D)={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (2.64)
In complete analogy with the subtraction (4.19), let us now introduce the subtraction for the YM part as follows:
$\Pi^{s(YM)}_{\rho\sigma}(q;D)$ $=$ $\Pi^{YM}_{\rho\sigma}(q;D)-\Pi^{YM}_{\rho\sigma}(0;D)$ (2.65)
$=$ $\Pi^{YM}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{YM}(D)$
where $\Pi^{s(YM)}_{\rho\sigma}(0;D)=0$, by definition, and
$\Delta^{2}_{YM}(D)=\Delta^{2}_{g}(D)+\Delta^{2}_{t}(D),$ (2.66)
as it follows from eq. (B.2).
Instead of the tensor decompositions (4.22), for the YM parts one gets
$\Pi^{s(YM)}_{\rho\sigma}(q)=T_{\rho\sigma}(q)q^{2}\Pi^{s(YM)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(YM)}_{l}(q^{2}),$ (2.67)
and
$\Pi^{YM}_{\rho\sigma}(q)=T_{\rho\sigma}(q)q^{2}\Pi^{YM}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{YM}_{l}(q^{2}).$ (2.68)
In all the quantities above and below the dependence on $D$ is omitted, for simplicity, and will be restored when necessary. Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary. However, all the invariant functions $\Pi^{s(YM)}_{t}(q^{2})$, $\Pi^{s(YM)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the subtraction (B.5).
Substituting the relations (B.7)-(B.8) into the relations (B.4)-(B.5), by omitting the tedious tensor algebra derivations, one finally obtains
$\Pi^{s(YM)}_{t}(q^{2})$ $=$ $\Pi^{YM}_{t}(q^{2})-{\Delta^{2}_{YM}(D)\over q^{2}},$
$\Pi^{s(YM)}_{l}(q^{2})$ $=$ $-{(\xi_{0}-\xi)\over\xi\xi_{0}}+{\Delta^{2}_{YM}(D)\over q^{2}}.$ (2.69)
However, the invariant function $\Pi^{s(YM)}_{l}(q^{2})$ cannot have the pole-type singularities, by definition, as underlined just above. So that one has to put the QD but regularized constant $\Delta^{2}_{YM}(D)=0$ on this mathematical ground. Taking into account that the quark constant $\Delta^{2}_{q}=0$ as well, one concludes that all the QD but regularized constants should be removed from theory, i.e., to be put zero as follows:
$\Delta^{2}_{q}=\Delta^{2}_{g}(D)=\Delta^{2}_{t}(D)=0,$ (2.70)
because of the relation (B.6) and all of them are independent from each other. If $\Delta^{2}_{t}(D)=0$ then from the relation (B.2) it follows that $\Pi^{YM}_{\rho\sigma}(q;D)=\Pi^{g}_{\rho\sigma}(q;D)$. But from the satisfied transverse relation $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$ and previous equality it follows that $q_{\rho}q_{\sigma}\Pi^{YM}_{\rho\sigma}(q;D)=0$ as well. Then from the relation (B.4) one gets $\xi_{0}=\xi$. The relations (B.10) and this equality determines the particular, or, equivalently, conventional QCD, investigated in section 6. Here it has been obtained by treating all the QD constants on the same footing, completely ignoring their principle difference from the physical point of view, which has been emphasized throughout this work.
This happened because in the previous derivations we deliberately have ignored that the subtraction (B.5), on account of the relation (B.2), in fact, looks like
$\Pi^{s(YM)}_{\rho\sigma}(q;D)=\Pi^{g}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{g}(D),$ (2.71)
i.e., $\Pi^{s(YM)}_{\rho\sigma}(q;D)=\Pi^{s(g)}_{\rho\sigma}(q;D)$, as it follows from eq. (4.18). Let us underline that the disappearance of the tadpole term is not a specific feature of our subtraction scheme. This always will happen in any subtraction scheme, since the constant minus the same constant is always zero, as pointed out in section 4. That is why explicitly present in the initial expressions it will remain in the final ones, compare eq. (2.5) with eq. (5.27), and thus with eq. (5.34).
Substituting further the tensor decomposition (B.7) and second of the decompositions (4.22) into the subtraction (B.11), and again omitting some tedious tensor algebra derivations, one obtains
$\Pi^{s(YM)}_{t}(q^{2})$ $=$ $\Pi^{g}_{t}(q^{2})-{\Delta^{2}_{g}(D)\over q^{2}},$
$\Pi^{s(YM)}_{l}(q^{2})$ $=$ $\Pi^{g}_{l}(q^{2})+{\Delta^{2}_{g}(D)\over q^{2}}.$ (2.72)
The general relation (B.4), on account of the first line of the subtraction (B.5), becomes
$q_{\rho}q_{\sigma}[\Pi^{s(YM)}_{\rho\sigma}(q;D)+\Pi^{YM}_{\rho\sigma}(0;D)]={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (2.73)
Substituting now the decomposition (B.7) and the relation $\Pi^{YM}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{YM}(D)$, which follows from the relations (B.5), into the previous equation, and omitting again tensor algebra derivations, one obtains
$\Pi^{s(YM)}_{l}(q^{2})=-{(\xi_{0}-\xi)\over\xi\xi_{0}}+{(\Delta^{2}_{g}(D)+\Delta^{2}_{t}(D))\over q^{2}}.$ (2.74)
Equating now this relation to the second of the relations (B.12), and doing simple algebra, one finally arrives at
${(\xi_{0}-\xi)\over\xi\xi_{0}}q^{2}=\Delta^{2}_{t}(D),$ (2.75)
which is nothing else but the general constraint (3.15). Note that in deriving this expression, we again use that from the satisfied transverse relation for the gluon part $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$ it follows that $\Pi^{g}_{l}(q^{2})=0$. As in the case of the conventional QCD, the properties of the gluon degrees of freedom have been used at the last stage.
Even introducing the YM contribution to the full gluon self-energy in eq. (B.1), which hides the tadpole term, but correctly treated its subtraction (B.5) by the subtraction (B.11), one again arrives at the exact constraint (B.15), nevertheless. Treated it literally (i.e., by eq. (B.5) only) leads to the conventional QCD. In this case the significant piece of the information on the true dynamical and gauge structures of the QCD vacuum is losing. In other words, the conventional QCD hides the true role of the tadpole term. Contrary, in general QCD, based on the exact solution of eq. (B.15), this information is not losing and the true role of the tadpole has been revealed.
Concluding, let us emphasize that the exact constraint (3.15) has been obtained from the very beginning before introducing any proper subtraction scheme. This underlines its general and unique character. At the same time, the clarifying algebraic derivations performed here, leading to the relations (B.10), heavily depend on the not accurate (i.e., literally) treatment of the subtraction (B.5), as explained just above.
Appendix C Renormalized ST identity
Substituting eq. (3.16) into the ST identity (3.7), one obtains
$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi=i{\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)}.$ (3.76)
Let us introduce the non-perturbative (NP) renormalization constant for the tadpole term as follows: $\Delta^{2}_{t}(D)=Z_{\Delta}\Delta^{2}$, where $\Delta^{2}$ is finite and positive, by definition, while the NP renormalization constant $Z_{\Delta}$ depends on all the unphysical parameters, such as regulating parameters, mentioned above, etc. In principle, the string tension in the linear rising potential between heavy quarks has to be somehow related to the finite $\Delta^{2}$, since it is connected to the transverse projection operator structure, as underlined above. It is well-known that the ansatz $D_{\mu\nu}(q)\sim\delta_{\mu\nu}\Delta^{2}/(q^{2})^{2}$ just leads to a such potential. After doing some algebra, the ST identity for the renormalized full gluon propagator becomes
$q_{\mu}q_{\nu}D^{R}_{\mu\nu}(q)=i\xi^{R}=i{\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}},$ (3.77)
where the renormalized gauge-fixing parameters are defined as follows: $\tilde{\xi}_{0}=Z_{\Delta}\xi_{0},\ \xi^{R}=Z_{\Delta}\xi$ and $D^{R}_{\mu\nu}(q)=Z_{\Delta}D_{\mu\nu}(q)$, i.e., the ST identity (C.1) is expressed in terms of the finite quantities only. The correct PT limit is maintained, of course.
Let us now demonstrate the interesting feature of the generalized gauge (C.2). It provides the self-consistency condition for the gauge choice in QCD as follows:
${\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}}={aq^{2}\over q^{2}+a\Delta^{2}},$ (3.78)
[…]
${\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}}={q^{2}\over\Delta^{2}},$ (3.79)
which is only satisfied at $q^{2}=0$, i.e., there is no any condition for the gauge-fixing parameter. The formulated self-consistency condition (C.3) points out on the inconsistency of the canonical gauge $\tilde{\xi}_{0}=\infty$ in QCD. Briefly, if everything is expressed in the terms of the finite quantities then the formal canonical gauge is forbidden to use in QCD, according to our approach to this theory.
Let us now apply the above-developed formalism to the renormalized massive gluon fields. For this purpose, it is instructive to re-write eq. (7.46) for the free massive gluon propagator in the terms of the renormalized quantities. But before it is convenient to introduce the following notations: $\tilde{\xi}_{0}=\tilde{\lambda}^{-1}$ and $M^{2}=M^{2}_{g}$, and the corresponding renormalization constant $Z_{\Delta}=1$ in this case. Then eq. (7.46) will look like
$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2}_{g})}\left[\delta_{\mu\nu}-{(1-\tilde{\lambda}^{-1})q_{\mu}q_{\nu}\over(q^{2}-\tilde{\lambda}^{-1}M^{2}_{g})}\right].$ (3.80)
From this observation it follows that our approach makes it possible for gluon to acquire mass dynamically without introducing into the theory some extra degrees of freedom such as Higgs fields [41], mentioned above, and $M^{2}_{g}$ can be treated as exactly defined gluon pole mass.
If one puts the expression (C.5) on the mass-shell $q^{2}=M^{2}_{g}$, then one arrives at
$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2}_{g})}\left[\delta_{\mu\nu}-{q_{\mu}q_{\nu}\over M^{2}_{g}}\right],$ (3.81)
which leads to the non-renormalizable behaviour in the PT $q^{2}\rightarrow\infty$ limit (i.e., it is constant and not $\sim 1/q^{2}$ as it is required for the renormalizable behaviour). However, the same expression (C.6) will be obtained if one expresses (C.5) in the canonical gauge ${\tilde{\lambda}}^{-1}=\infty$. As explained above, this gauge is not to be used in QCD. This means that within the generalized gauge-fixing function approach the free massive gluons cannot be put on the mass-shell. The finite values for the gauge-fixing parameter ${\tilde{\lambda}}^{-1}$ are only permitted in order to guarantee the theory to be renormalizable.The expression (C.5) has the correct PT limit, being off-mass-shell object at any finite ${\tilde{\lambda}}^{-1}$. So that, the canonical gauge is rather formal than real within the generalized gauge-fixing function approach to QCD. The canonical gauge and its connection to the structure of the full gluon propagator at the mass-shell will be investigated in our future work, where the NP renormalization theory of the massive gluon fields will be present in detail.
Appendix D Preservation of the gauge symmetry in the QED ground state
The propagation of the photon in the QED vacuum is formally described by eq. (2.1). Contrary to QCD (compare with eq. (2.2), the difference is that there is only one contribution to the full photon self-energy comes from the vacuum polarization tensor. Its skeleton term is similar present in Fig. 1 by the first diagram. Analytically it looks like
$\Pi^{ph}_{\rho\sigma}(q)\sim\int d^{4}l\gamma_{\rho}S(l-q)\Gamma_{\sigma}(l,q)S(l),$ (4.82)
where all the overall numerical factors are omitted, for simplicity. Here $S$ and $G$ represent the full electron propagator and the full electron-photon vertex, respectively. The full photon and its free counterpart are analytically present similar to eq. (3.8) and eq. (3.9), respectively, as well as their ST identities by relations (3.7) and (3.10).
Similar to the relation (3.11) for the transverse relation in this case, one obtains
$q_{\rho}q_{\sigma}\Pi^{ph}_{\rho\sigma}(q;D)={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (4.83)
It is well-known that because of the electric charge conversation, the transverse relation for the vacuum polarization tensor in QED is always satisfied, i.e., it is zero. Then from the previous equation immediately follows
$q_{\rho}q_{\sigma}\Pi^{ph}_{\rho\sigma}(q;D)=0,\quad\rightarrow\quad\xi_{0}=\xi.$ (4.84)
Let us now introduce the corresponding subtraction scheme similar to the relation (4.19) and their independent tensor decompositions (4.22), namely
$\Pi^{s(ph)}_{\rho\sigma}(q)=\Pi^{ph}_{\rho\sigma}(q)-\Pi^{ph}_{\rho\sigma}(0)=\Pi^{ph}_{\rho\sigma}(q)-\delta_{\rho\sigma}\Delta^{2}_{ph},$ (4.85)
where $\Pi^{s(ph)}_{\rho\sigma}(0)=0$, by definition, and
$\Pi^{s(ph)}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(ph)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(ph)}_{l}(q^{2}),$
$\Pi^{ph}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{ph}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{ph}_{l}(q^{2}),$ (4.86)
respectively. Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary. However, the both invariant functions $\Pi^{s(ph)}_{t}(q^{2})$, $\Pi^{s(ph)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the initial subtraction (D.4).
Substituting all these decompositions (D.5) into the subtraction (D.4), and doing some tensor algebra derivations, one finally obtains
$\Pi^{s(ph)}_{t}(q^{2})$ $=$ $\Pi^{ph}_{t}(q^{2})-{\Delta^{2}_{ph}\over q^{2}},$
$\Pi^{s(ph)}_{l}(q^{2})$ $=$ $\Pi^{ph}_{l}(q^{2})+{\Delta^{2}_{ph}\over q^{2}}.$ (4.87)
From the first of the relations (D.3) and the second of the relations (D.5) it follows that $\Pi^{ph}_{l}(q^{2})=0$. Then the second of the relations (D.6) will be reduced to
$\Pi^{s(ph)}_{l}(q^{2})={\Delta^{2}_{ph}\over q^{2}},$ (4.88)
which is impossible since the corresponding invariant function can not have the pole-type singularities, by definition, as explained above. One has to put the QD but regularized constant to zero on this mathematical basis, i.e., $\Delta^{2}_{ph}=0$, and thus $\Pi^{s(ph)}_{l}(q^{2})=0$ as well.
As we already know, the both relations
$\xi=\xi_{0},\quad\Delta^{2}_{ph}=0$ (4.89)
are necessary conditions for the coincidence of the gauge symmetry of the QED Lagrangian with the gauge symmetry of its ground state (like it occurs for the conventional QCD). This is our proof of the Elitzur theorem [42] for the continuos abelian fields within our approach, based on the formalism of the dynamical equations of motion for the gauge particle Euclidean Green’s functions.